### v4

# Pfister’s local-global principle for Azumaya algebras with involution

Abstract.
We prove Pfister’s local-global principle for hermitian forms over Azumaya algebras with involution over semilocal rings, and show in particular that the Witt group of nonsingular hermitian forms is $2$-primary torsion. Our proof relies on a hermitian version of Sylvester’s law of inertia, which is obtained from an investigation of the connections between a pairing of hermitian forms extensively studied by Garrel, signatures of hermitian forms, and positive semidefinite quadratic forms.
Key words and phrases:
Azumaya algebras, involutions, hermitian forms, Witt groups, torsion, positivity
2020 Mathematics Subject Classification:
16H05, 11E39, 13J30, 16W10
Introduction
Pfister’s local-global principle is a fundamental result in the algebraic theory of quadratic forms over fields. It states that the torsion in the Witt ring is $2$-primary, and that a nonsingular quadratic form represents a torsion element in this ring if and only if its signature (the difference between the number of positive and the number of negative entries in any diagonalization of the form according to Sylvester’s “law of inertia”) is zero at all orderings of the field. The above facts are of course well-known, and can easily be found in standard references such as [28] or [19].
The main result of this paper (Theorem 5.6) is Pfister’s local-global principle for nonsingular hermitian forms over Azumaya algebras with involution over semilocal commutative rings in which $2$ is a unit. (The special case of central simple algebras with involution was treated in [21] and [27]; see also [7].) The assumption that the base ring is semilocal is minimal in the sense that Pfister’s local-global principle is known to hold for nonsingular quadratic forms over such rings, but not in general. We refer to [5] for more details.
Our version of Sylvester’s law of inertia (Theorem 5.1) is used in the proof of the main result, which is inspired by Marshall’s proof of Pfister’s local-global principle in the context of abstract Witt rings, cf. [23]. A crucial ingredient in our proof is a certain pairing of forms investigated by Garrel in his 2023 paper [10]. We were also very fortunate that we could put many results from the recent papers by First [8] and Bayer-Fluckiger, First and Parimala [6] to good use.
1. Preliminaries
In this paper all rings are assumed unital and associative with $2$ invertible. We identify quadratic forms over commutative rings with symmetric bilinear forms, and assume that all fields are of characteristic different from $2$. Rings are not assumed to be commutative unless explicitly indicated. Our main references for rings with involution and hermitian forms are [16] and [8].
1.1. Hermitian forms over rings with involution
Let $(A,\sigma)$ be a ring with involution and let $\varepsilon\in Z(A)$ be such that $\varepsilon\sigma(\varepsilon)=1$. We denote the category of $\varepsilon$-hermitian modules over $(A,\sigma)$ by $\mathfrak{Herm}^{\varepsilon}(A,\sigma)$. The objects of $\mathfrak{Herm}^{\varepsilon}(A,\sigma)$ are pairs $(M,h)$, where $M$ is a finitely generated projective right $A$-module and $h\colon M\times M\to A$ is an $\varepsilon$-hermitian form. Since we always assume that $2\in A^{\times}$, all hermitian modules are even. We denote the category of nonsingular $\varepsilon$-hermitian modules over $(A,\sigma)$ (also known as $\varepsilon$-hermitian spaces) by $\mathfrak{H}^{\varepsilon}(A,\sigma)$. The morphisms of $\mathfrak{Herm}^{\varepsilon}(A,\sigma)$ and $\mathfrak{H}^{\varepsilon}(A,\sigma)$ are the isometries. We denote isometry by $\simeq$. See [16, I, Sections 2 and 3] for more details. If $\varepsilon=1$, we simply write $\mathfrak{Herm}(A,\sigma)$ and $\mathfrak{H}(A,\sigma)$. It is common to say hermitian form instead of hermitian module.
We denote the Witt group of nonsingular $\varepsilon$-hermitian forms over $(A,\sigma)$ by $W^{\varepsilon}(A,\sigma)$ and note that since $2\in A^{\times}$, metabolic forms are hyperbolic, cf. [8, Section 2.2] for a succinct presentation.
For $a\in A^{\times}$, we denote the inner automorphism $A\to A,\ x\mapsto axa^{-1}$ by $\operatorname{Int}(a)$. We define $\operatorname{Sym}^{\varepsilon}(A,\sigma):=\{x\in A\mid\sigma(x)=\varepsilon x\}$ and $\operatorname{Sym}^{\varepsilon}(A^{\times},\sigma):=\operatorname{Sym}^{\varepsilon}(A,\sigma)\cap A^{\times}$. We also write $\operatorname{Sym}(A,\sigma)$ instead of $\operatorname{Sym}^{1}(A,\sigma)$ and $\operatorname{Skew}(A,\sigma)$ instead of $\operatorname{Sym}^{-1}(A,\sigma)$. For $a_{1},\ldots,a_{\ell}\in\operatorname{Sym}^{\varepsilon}(A,\sigma)$ we denote by $\langle a_{1},\ldots,a_{\ell}\rangle_{\sigma}$ the diagonal $\varepsilon$-hermitian form
$A^{\ell}\times A^{\ell}\to A,\ (x,y)\mapsto\sum_{i=1}^{\ell}\sigma(x_{i})a_{i}y_{i}.$
Let $(M,h)\in\mathfrak{Herm}^{\varepsilon}(A,\sigma)$. We denote by $D_{(A,\sigma)}(h):=\{h(x,x)\mid x\in M\}$ the set of elements of $A$ represented by $h$.
If $(M,h)\in\mathfrak{H}^{\varepsilon}(A,\sigma)$, the adjoint involution of $h$ is the involution $\operatorname{ad}_{h}$ on the ring $\operatorname{End}_{A}(M)$ implicitly defined by
(1.1) $h(x,\operatorname{ad}_{h}(f)(y))=h(f(x),y)$
for all $x,y\in M$ and all $f\in\operatorname{End}_{A}(M)$, cf. [8, Section 2.4].
Consider a second ring with involution $(B,\tau)$ and let $(S,\iota)$ be a commutative ring with involution such that $(A,\sigma)$ and $(B,\tau)$ are $(S,\iota)$-algebras with involution in the sense of [16, I, (1.1)], i.e., $A$ and $B$ are both $S$-algebras and the involutions $\sigma$ and $\tau$ are compatible with $\iota$:
$\sigma(sa)=\iota(s)\sigma(a),\ \tau(sb)=\iota(s)\tau(b),\quad\forall a\in A,b\in B,s\in S.$
Then $(A\otimes_{S}B,\sigma\otimes\tau)$ is an $(S,\iota)$-algebra with involution and, if $(M_{1},h_{1})\in\mathfrak{Herm}^{\varepsilon_{1}}(A,\sigma)$ and $(M_{2},h_{2})\in\mathfrak{Herm}^{\varepsilon_{2}}(B,\tau)$, then
$(M_{1}\otimes_{S}M_{2},h_{1}\otimes h_{2})\in\mathfrak{Herm}^{\varepsilon_{1}\varepsilon_{2}}(A\otimes_{S}B,\sigma\otimes\tau).$
If $(M_{1},h_{1})\simeq(M_{1}^{\prime},h_{1}^{\prime})$ in $\mathfrak{Herm}^{\varepsilon_{1}}(A,\sigma)$ and $(M_{2},h_{2})\simeq(M_{2}^{\prime},h_{2}^{\prime})$ in $\mathfrak{Herm}^{\varepsilon_{2}}(B,\tau)$, then
$(M_{1}\otimes_{S}M_{2},h_{1}\otimes h_{2})\simeq(M_{1}^{\prime}\otimes_{S}M_{2}^{\prime},h_{1}^{\prime}\otimes h_{2}^{\prime}).$
Let $(M,h)\in\mathfrak{Herm}^{\varepsilon}(A,\sigma)$ and $(N,\varphi)\in\mathfrak{Herm}^{\mu}(S,\iota)$. Since $(A\otimes_{S}S,\sigma\otimes\iota)\cong(A,\sigma)\cong(S\otimes_{S}A,\iota\otimes\sigma)$ as rings with involution, it is not difficult to see that upon identifying $A\otimes_{S}S$ with $A$,
$h\otimes_{S}\varphi(m_{1}\otimes n_{1},m_{2}\otimes n_{2})=h(m_{1},m_{2})\varphi(n_{1},n_{2})$
for all $m_{1},m_{2}\in M$ and $n_{1},n_{2}\in N$, and
$(M\otimes_{S}N,h\otimes_{S}\varphi)\simeq(N\otimes_{S}M,\varphi\otimes_{S}h)$
in $\mathfrak{Herm}^{\varepsilon\mu}(A,\sigma)$.
Lemma 1.1.
Let $R$ be a commutative ring, assume that $A$ is an $R$-algebra and that $\sigma$ is an $R$-linear involution on $A$. Let $\iota=\sigma|_{Z(A)}$. Let $u_{1},\ldots,u_{k}\in R$ and $(M,h)\in\mathfrak{Herm}^{\varepsilon}(A,\sigma)$. Then
$\langle u_{1},\ldots,u_{k}\rangle_{\iota}\otimes_{Z(A)}h\simeq\langle u_{1},\ldots,u_{k}\rangle\otimes_{R}h$
(under the canonical identifications $Z(A)\otimes_{Z(A)}A\cong R\otimes_{R}A\cong A$).
Proof.
It suffices to show that $\langle u\rangle_{\iota}\otimes_{Z(A)}h\simeq\langle u\rangle\otimes_{R}h$ for $u\in R$. This follows from the isometries $\langle u\rangle_{\iota}\simeq\langle u\rangle\otimes_{R}\langle 1\rangle_{\iota}$, $\langle 1\rangle_{\iota}\otimes_{Z(A)}h\simeq h$ (which are straightforward, using the observations preceding the lemma) and associativity of the tensor product. ∎
We finish this section with a well-known result for which we could not find a reference:
Lemma 1.2.
Let $R$ be a commutative ring, assume that $A$ and $B$ are $R$-algebras and that $\sigma$ and $\tau$ are $R$-linear involutions on $A$ and $B$, respectively. Let $(M,\varphi)\in\mathfrak{H}^{\varepsilon_{1}}(A,\sigma)$ and $(N,\psi)\in\mathfrak{H}^{\varepsilon_{2}}(B,\tau)$. Then the map
$\xi\colon\operatorname{End}_{A}(M)\otimes_{R}\operatorname{End}_{B}(N)\to\operatorname{End}_{A\otimes_{R}B}(M\otimes_{R}N)$
induced by $\xi(f\otimes g)=[x\otimes y\mapsto f(x)\otimes g(y)]$ yields an isomorphism of $R$-algebras with involution
$(\operatorname{End}_{A}(M)\otimes_{R}\operatorname{End}_{B}(N),\operatorname{ad}_{\varphi}\otimes\operatorname{ad}_{\psi})\cong(\operatorname{End}_{A\otimes_{R}B}(M\otimes_{R}N),\operatorname{ad}_{\varphi\otimes\psi}).$
Proof.
The map $\xi$ is an isomorphism of $R$-algebras by [9, Theorem 1.3.26 and Corollary 1.3.27]. To finish the proof it suffices to check that $\operatorname{ad}_{\varphi\otimes\psi}(\xi(f\otimes g))=\xi(\operatorname{ad}_{\varphi}(f)\otimes\operatorname{ad}_{\psi}(g))$ using the definition of adjoint involution (1.1), which is a straightforward computation. ∎
1.2. Quadratic étale algebras
Let $R$ be a commutative ring and let $S$ be a quadratic étale $R$-algebra. We recall some results from [8, Section 1.3] and [16, I, (1.3.6)]. The algebra $S$ has a unique standard involution $\vartheta$, and the trace $\operatorname{Tr}_{S/R}$ satisfies
(1.2) $\operatorname{Tr}_{S/R}(x)=\vartheta(x)+x\quad\text{for all $x\in S$.}$
Furthermore, $\operatorname{Tr}_{S/R}$ is an involution trace for $\vartheta$ (cf. [16, I, Proposition 7.3.6]) and thus if $h\in\mathfrak{Herm}(S,\vartheta)$ is nonsingular, then $\operatorname{Tr}_{S/R}(h)$ is nonsingular by [16, I, Proposition 7.2.4]. Furthermore, if $h$ is hyperbolic, then so is the quadratic form $\operatorname{Tr}_{S/R}(h)$ by the first paragraph of [16, p. 41]. The converse holds if $R$ is semilocal by [8, Corollary 8.3].
If $R$ is connected and $S$ is not connected, then $S\cong R\times R$ as $R$-algebras, and $\vartheta\colon(x,y)\mapsto(y,x)$ is the exchange involution.
If $R$ is semilocal, then there exists $\lambda\in S$ such that $\lambda^{2}\in R^{\times}$, $\vartheta(\lambda)=-\lambda$, and $\{1,\lambda\}$ is an $R$-basis of $S$, cf. [8, Lemma 1.19].
1.3. Azumaya algebras with involution
Let $R$ be a commutative ring. Recall from [16, III, (5.1)] that an $R$-algebra $A$ is an Azumaya $R$-algebra if $A$ is a faithful finitely generated projective $R$-module and the sandwich map
(1.3) $\mathrm{sw}\colon A\otimes_{R}A^{\mathrm{op}}\to\operatorname{End}_{R}(A),\ a\otimes b^{\mathrm{op}}\mapsto[x\mapsto axb]$
is an isomorphism of $R$-algebras. Here $A^{\mathrm{op}}$ denotes the opposite algebra of $A$, which coincides with $A$ as an $R$-module, but with twisted multiplication $a^{\mathrm{op}}b^{\mathrm{op}}=(ba)^{\mathrm{op}}$. It is clear that $A^{\mathrm{op}}$ is also an Azumaya $R$-algebra.
The centre $Z(A)$ is equal to $R$ and $\operatorname{End}_{R}(A)$ is again an Azumaya $R$-algebra. More generally, if $M$ is a faithful finitely generated projective right $R$-module, then $\operatorname{End}_{R}(M)$ is an Azumaya $R$-algebra. If $A$ and $B$ are Azumaya $R$-algebras, their tensor product $A\otimes_{R}B$ is again an Azumaya $R$-algebra.
First’s paper [8] contains a wealth of information about Azumaya algebras, with and without involution, and we refer to it for a number of definitions and results that we recall in the remainder of this section.
Proposition 1.3 ([8, Proposition 1.1]).
$A$ is Azumaya over $Z(A)$ and $Z(A)$ is finite étale over $R$ if and only if $A$ is projective as an $R$-module and separable as an $R$-algebra.
Since the behaviour of the involution on the centre plays an important role in the study of algebras with involution, this result helps motivate the following:
Definition 1.4 ([8, Section 1.4]).
We say that $(A,\sigma)$ is an Azumaya algebra with involution over $R$ if the following conditions hold:
  • $A$ is an $R$-algebra with $R$-linear involution $\sigma$;
  • $A$ is separable projective over $R$;
  • the homomorphism $R\rightarrow A$, $r\mapsto r\cdot 1_{A}$ identifies $R$ with $\operatorname{Sym}(Z(A),\sigma)$.
Let $(A,\sigma)$ be an Azumaya algebra with involution over $R$. Note that $A$ is Azumaya over $Z(A)$ by Proposition 1.3, but may not be Azumaya over $R$. Indeed, “Azumaya algebra with involution” means “Azumaya algebra-with-involution” rather than “Azumaya-algebra with involution”.
The following lemma makes the connection between Definition 1.4 and a different definition of Azumaya algebra with involution (the first sentence of Lemma 1.5) that is introduced in [25, Section 4]:
Lemma 1.5.
Let $A$ be an $R$-algebra with $R$-linear involution such that $A$ is an Azumaya algebra over $Z(A)$, $Z(A)$ is $R$ or a quadratic étale extension of $R$, and $R=\operatorname{Sym}(Z(A),\sigma)$. Then $(A,\sigma)$ is an Azumaya algebra with involution over $R$.
The converse holds if $R$ is connected.
Proof.
The first statement is a direct consequence of Proposition 1.3. For the converse, assume that $R$ is connected and that $(A,\sigma)$ is an Azumaya algebra with involution over $R$. Then $A$ is Azumaya over $Z(A)$ by Proposition 1.3, and $Z(A)$ is $R$ or a quadratic étale extension of $R$ by [8, Proposition 1.21]. ∎
Remark 1.6.
If $R=F$ is actually a field, then $(A,\sigma)$ is an Azumaya algebra with involution over $R$ if and only if it is a central simple $F$-algebra with involution in the sense of [17, 2.A, 2.B].
We recall the following, proved in [8, second paragraph of Section 1.4]:
Proposition 1.7 (Change of base ring).
Let $T$ be a commutative $R$-algebra. Then $Z(A\otimes_{R}T)=Z(A)\otimes_{R}T$ and $(A\otimes_{R}T,\sigma\otimes\mathrm{id})$ is an Azumaya algebra with involution over $T$ (and in particular $Z(A\otimes_{R}T)\cap\operatorname{Sym}(A\otimes_{R}T,\sigma\otimes\operatorname{id})=T$).
Corollary 1.8.
Let $T$ be a commutative $R$-algebra which is also a field. Then either $Z(A\otimes_{R}T)=T$ or $Z(A\otimes_{R}T)$ is a quadratic étale extension of $T$, and $(A\otimes_{R}T,\sigma\otimes\mathrm{id})$ is a central simple $T$-algebra with involution.
Remark 1.9.
Let $S$ be a quadratic étale $R$-algebra with standard involution $\vartheta$. If $T$ is a commutative $R$-algebra, then $S\otimes_{R}T$ is a quadratic étale $T$-algebra, cf. [9, Propositions 2.3.4 and 9.2.5].
Moreover, $(S,\vartheta)$ is an Azumaya algebra with involution over $R$ in the sense of [25, Section 4] (i.e., it has the properties listed in the first sentence of Lemma 1.5), and therefore an Azumaya algebra with involution over $R$ in the sense of [8] (i.e., in the sense of this paper), by Lemma 1.5. Therefore $T=\operatorname{Sym}(S\otimes_{R}T,\vartheta\otimes\operatorname{id})$ by Proposition 1.7, and it follows that $\vartheta\otimes\operatorname{id}$ is the standard involution of $S\otimes_{R}T$. Then (1.2) applies and yields, for all $x\in S$,
$\operatorname{Tr}_{S\otimes_{R}T/T}(x\otimes 1)=\operatorname{Tr}_{S/R}(x)\otimes 1.$
We collect some results about ideals of $A$:
Proposition 1.10.
The following properties hold:
  1. (1)
    If $I$ is an ideal of $R$, then $A\otimes_{R}(R/I)\cong A/AI$;
  2. (2)
    $J(A)=A\cdot J(R)$, where $J$ denotes the Jacobson radical;
  3. (3)
    If $a\in A$ is such that $a\in(A/A\mathfrak{m})^{\times}$ for every maximal ideal $\mathfrak{m}$ of $R$, then $a\in A^{\times}$.
Proof.
(1) This is well-known. (2) Is [8, Lemma 1.5], since $A$ is a separable projective $R$-algebra. (3) Is [8, Lemma 1.6], using item (1). ∎
Definition 1.11 (See [8, Section 1.4]).
We say that the involution $\sigma$ on $A$ is of orthogonal, symplectic or unitary type at $\mathfrak{p}\in\operatorname{Spec}R$ if $(A\otimes_{R}\mathrm{qf}(R/\mathfrak{p}),\sigma\otimes\operatorname{id})$ is a central simple algebra with involution of orthogonal, symplectic or unitary type, respectively, over the quotient field $\mathrm{qf}(R/\mathfrak{p})$, and that $\sigma$ is of orthogonal, symplectic or unitary type if it is of orthogonal, symplectic or unitary type, respectively, at all $\mathfrak{p}\in\operatorname{Spec}R$.
We often simply say that $\sigma$ is orthogonal, symplectic or unitary (at $\mathfrak{p}$). We refer to [17] for the definitions of orthogonal, symplectic and unitary involutions on central simple algebras.
Proposition 1.12 (See [8, Proposition 1.21] for a more detailed statement).
Assume that $R$ is connected. Then precisely one of the following holds:
  1. (1)
    $\sigma$ is of orthogonal type, $Z(A)=R$, and $\sigma|_{Z(A)}=\mathrm{id}_{Z(A)}$;
  2. (2)
    $\sigma$ is of symplectic type, $Z(A)=R$, and $\sigma|_{Z(A)}=\mathrm{id}_{Z(A)}$;
  3. (3)
    $\sigma$ is of unitary type, $Z(A)$ is a quadratic étale $R$-algebra, and $\sigma|_{Z(A)}$ is the standard involution on $Z(A)$.
Remark 1.13.
If $(A,\sigma)$ is a central simple algebra over a field, the involution $\sigma$ is called of the first kind if $\sigma|_{Z(A)}=\operatorname{id}_{Z(A)}$ and of the second kind otherwise (so that “second kind” is the same as “unitary type”), cf. [17].
Lemma 1.14.
Let $R$ be semilocal connected, let $T$ be a quadratic étale $R$-algebra, and let $\tau$ be an $R$-linear involution on $T$. If $T$ is connected, then every nonsingular hermitian form over $(T,\tau)$ is diagonalizable.
Proof.
The algebra $T$ is semilocal by [16, VI, Proposition 1.1.1]. Since $T$ is also connected, every projective $T$-module is free by [11]. Furthermore, $2\in T^{\times}$.
If $\tau=\operatorname{id}$, the result follows since every quadratic form over $T$ is diagonalizable (see [5, Proposition 3.4]).
If $\tau\not=\operatorname{id}$. Then $\tau$ is the standard involution on $T$ by [8, Lemma 1.17], and $(T,\tau)$ is an Azumaya algebra with involution over $R$, cf. Remark 1.9. Since $Z(T)\not=R$, and in particular $\tau$ is not symplectic, we may conclude by [8, Proposition 2.13]. ∎
Theorem 1.15 (Hermitian Morita equivalence).
Let $R$ be semilocal connected and let $(B,\tau)$ be an Azumaya algebra with involution over $R$ such that $A$ and $B$ are Brauer equivalent and $\sigma|_{S}=\tau|_{S}$, where $S:=Z(A)=Z(B)$. Assume that $S$ is connected. Then there exists $\delta\in\{-1,1\}$ such that the categories $\mathfrak{Herm}^{\varepsilon}(A,\sigma)$ and $\mathfrak{Herm}^{\delta\varepsilon}(B,\tau)$ are equivalent, and this equivalence induces an isomorphism of Witt groups $W^{\varepsilon}(A,\sigma)\cong W^{\delta\varepsilon}(B,\tau)$. Specifically, if $\sigma|_{S}=\operatorname{id}_{S}$, then $\delta=1$ if $\sigma$ and $\tau$ are both orthogonal or both symplectic and $\delta=-1$ otherwise; if $\sigma|_{S}\not=\operatorname{id}_{S}$, then $\delta$ can be chosen freely in $\{-1,1\}$.
Proof.
If $\sigma|_{S}\not=\operatorname{id}_{S}$, then $S$ is a quadratic étale $R$-algebra by Proposition 1.12 and is thus semilocal by [16, VI, Proposition 1.1.1].
Since $A$ and $B$ are Brauer equivalent, there exist faithful finitely generated projective $S$-modules $P$ and $Q$ such that $A\otimes_{S}\operatorname{End}_{S}(P)\cong B\otimes_{S}\operatorname{End}_{S}(Q)$, cf. [16, III, (5.3)]. Since $S$ is semilocal connected, $P$ and $Q$ are free by [11]. It follows that there exist $k,\ell\in\mathbb{N}$ such that
(1.4) $\operatorname{End}_{A}(A^{k})\cong\operatorname{End}_{B}(B^{\ell}).$
Consider the nonsingular hermitian forms $(A^{k},\varphi:=k\times\langle 1\rangle_{\sigma})$ and $(B^{\ell},\psi:=\ell\times\langle 1\rangle_{\tau})$ over $(A,\sigma)$ and $(B,\tau)$, respectively. Since $\varphi$ and $\psi$ are hermitian, the involutions $\sigma$ and $\operatorname{ad}_{\varphi}$ are of the same type and the same is true for the involutions $\tau$ and $\operatorname{ad}_{\psi}$, cf. [8, Proposition 2.11]. In particular, $\sigma|_{S}=\operatorname{ad}_{\varphi}|_{S}$ and $\tau|_{S}=\operatorname{ad}_{\psi}|_{S}$.
The categories $\mathfrak{Herm}^{\varepsilon}(\operatorname{End}_{A}(A^{k}),\operatorname{ad}_{\varphi})$ and $\mathfrak{Herm}^{\varepsilon}(A,\sigma)$ are equivalent by hermitian Morita theory. The same is true for the categories $\mathfrak{Herm}^{\varepsilon}(\operatorname{End}_{B}(B^{\ell}),\operatorname{ad}_{\psi})$ and $\mathfrak{Herm}^{\varepsilon}(B,\tau)$. Furthermore, these equivalences respect orthogonal sums and send hyperbolic spaces to hyperbolic spaces, cf. [16, I, Theorem 9.3.5].
The involution $\operatorname{ad}_{\varphi}$ induces an involution $\omega$ on $\operatorname{End}_{B}(B^{\ell})$ of the same type via the isomorphism (1.4), and thus $\omega|_{S}=\operatorname{ad}_{\psi}|_{S}$ since $\sigma|_{S}=\tau|_{S}$. By the Skolem-Noether theorem [6, Theorem 8.6], $\omega$ differs from $\operatorname{ad}_{\psi}$ by an inner automorphism: there exists $\delta\in\{-1,1\}$ and a unit $u\in\operatorname{End}_{B}(B^{\ell})$ with $\operatorname{ad}_{\psi}(u)=\delta u$ such that $\omega=\operatorname{Int}(u)\circ\operatorname{ad}_{\psi}$, where $\delta$ can be freely chosen in $\{-1,1\}$ if $\sigma|_{S}\not=\operatorname{id}_{S}$. Moreover, $\delta$ is unique if $\sigma|_{S}=\operatorname{id}_{S}(=\operatorname{ad}_{\psi}|_{S})$: if $\operatorname{Int}(u)=\operatorname{Int}(u^{\prime})$, then $u=u^{\prime}s$ for some $s\in S^{\times}$, and it follows that $\operatorname{ad}_{\psi}(u)=\delta u$ if and only if $\operatorname{ad}_{\psi}(u^{\prime})=\delta u^{\prime}$.
By [16, I, (5.8)] (see also [8, Section 2.7]) there is an equivalence between the categories $\mathfrak{Herm}^{\varepsilon}(\operatorname{End}_{B}(B^{\ell}),\omega)$ and $\mathfrak{Herm}^{\delta\varepsilon}(\operatorname{End}_{B}(B^{\ell}),\operatorname{ad}_{\psi})$ (and an induced equivalence between the categories $\mathfrak{H}^{\varepsilon}(\operatorname{End}_{B}(B^{\ell}),\omega)$ and $\mathfrak{H}^{\delta\varepsilon}(\operatorname{End}_{B}(B^{\ell}),\operatorname{ad}_{\psi})$) which respects orthogonal sums and hyperbolic spaces. The equivalence of the categories $\mathfrak{Herm}^{\varepsilon}(A,\sigma)$ and $\mathfrak{Herm}^{\delta\varepsilon}(B,\tau)$ and the isomorphism of the Witt groups $W^{\varepsilon}(A,\sigma)$ and $W^{\delta\varepsilon}(B,\tau)$ follows.
It remains to show the claim about $\delta$ when $\sigma|_{S}=\operatorname{id}_{S}$. Since $\delta$ is unique, it suffices to check its value at any $\mathfrak{p}\in\operatorname{Spec}R$ (i.e., after tensoring over $R$ by $\mathrm{qf}(R/\mathfrak{p})$). By [17, Proposition 2.7] we know that $\delta=1$ if $\sigma$ and $\tau$ are both orthogonal or both symplectic at $\mathfrak{p}$ and $\delta=-1$ otherwise. The statement then follows, using Proposition 1.12. ∎
Remark 1.16.
Let $R$ be connected and let $(A,\sigma)$ be an Azumaya algebra with involution over $R$. Note that when $Z(A)$ is not connected, then $Z(A)\cong R\times R$ by Section 1.2 and Lemma 1.5. Hence, $W^{\varepsilon}(A,\sigma)=0$, cf. [8, Example 2.4].
1.4. Orderings, Spaces of signatures
Let $R$ be a commutative ring, and let $(A,\sigma)$ be an Azumaya algebra with involution over $R$. The real spectrum of $R$, $\operatorname{Sper}R$, is the set of all orderings on $R$, cf. [15, Definitions 3.3.1(a) and 3.3.4, and Proposition 3.3.5]. It is equipped with the Harrison topology, with subbasis given by the sets of the form
$\mathring{H}(r):=\{\alpha\in\operatorname{Sper}R\mid r>0\text{ at }\alpha\}=\{\alpha\in\operatorname{Sper}R\mid r\in\alpha\setminus-\alpha\},$
for all $r\in R$.
Let $\alpha\in\operatorname{Sper}R$. We often write $x\geq_{\alpha}y$ for $x-y\in\alpha$. The support of $\alpha$ is the prime ideal $\operatorname{Supp}(\alpha):=\alpha\cap-\alpha\in\operatorname{Spec}R$, and we denote by $\bar{\alpha}$ the ordering induced by $\alpha$ on the quotient field $\kappa(\alpha):=\mathrm{qf}(R/\operatorname{Supp}(\alpha))$, and by $k(\alpha)$ a real closure of $\kappa(\alpha)$ at $\bar{\alpha}$. Observe that $\kappa(\bar{\alpha}):=\mathrm{qf}(\kappa(\alpha)/\operatorname{Supp}(\bar{\alpha}))=\mathrm{qf}(\kappa(\alpha)/\{0\})=\kappa(\alpha)$, and that we can take $k(\bar{\alpha})=k(\alpha)$.
We also define
$(A(\alpha),\sigma(\alpha)):=(A\otimes_{R}\kappa(\alpha),\sigma\otimes\operatorname{id}),$
which is a central simple $\kappa(\alpha)$-algebra with involution by Corollary 1.8.
Remark 1.17.
Let $M$ be an $R$-module and let $\alpha\in\operatorname{Sper}R$. Then, if $x\in M\otimes_{R}\kappa(\alpha)$, there are $t\in\mathbb{N}$, $m_{i}\in M$, $s_{i}\in R$ and $r_{i}\in R\setminus\operatorname{Supp}(\alpha)$ such that
$x=\sum_{i=1}^{t}m_{i}\otimes\dfrac{\bar{s}_{i}}{\bar{r}_{i}}=\sum_{i=1}^{t}m_{i}\otimes\dfrac{\bar{s}^{\prime}_{i}}{\bar{r}^{\prime}}=\Bigl{(}\sum_{i=1}^{t}m_{i}s^{\prime}_{i}\Bigr{)}\otimes\dfrac{1}{\bar{r}^{\prime}},$
for some $s^{\prime}_{i}\in R$, $r^{\prime}\in R\setminus\operatorname{Supp}(\alpha)$. Therefore
$M\otimes_{R}\kappa(\alpha)=\Big{\{}m\otimes\dfrac{1}{\bar{r}}\,\Big{|}\,m\in M,\ r\in R\setminus\operatorname{Supp}(\alpha)\Big{\}}.$
We denote by $\operatorname{Sign}R$ the set of signatures of $R$, i.e., the space of all morphisms of rings from the Witt ring $W(R)$ to $\mathbb{Z}$. We recall some facts from [14, Section 5 up to p. 89]:
  • The set $\operatorname{Sign}R$ is equipped with the coarsest topology that makes all maps
    $\operatorname{sign}_{\bullet}\varphi\colon\operatorname{Sper}R\rightarrow\mathbb{Z},\ \alpha\mapsto\operatorname{sign}_{\alpha}\varphi$
    continuous, for $\varphi\in W(R)$. When $R$ is semilocal, a basis for this topology is given by the sets
    $H(u_{1},\ldots,u_{k}):=\{\tau\in\operatorname{Sign}R\mid\tau(u_{1})=\cdots=\tau(u_{k})=1\},$
    for $k\in\mathbb{N}$ and $u_{1},\ldots,u_{k}\in R^{\times}$.
  • Denoting by $\operatorname{Sper}^{\mathrm{max}}R$ the space of all elements of $\operatorname{Sper}R$ that are maximal for inclusion (equipped with the induced topology), the natural map
    $\operatorname{Sper}^{\mathrm{max}}R\rightarrow\operatorname{Sign}R,\ \alpha\mapsto\operatorname{sign}_{\alpha},$
    is continuous and surjective. If $R$ is semilocal, this map is a homeomorphism and we identify $\operatorname{Sign}R$ and $\operatorname{Sper}^{\mathrm{max}}R$. In general, we have continuous maps
    $\operatorname{Sper}^{\mathrm{max}}R\subseteq\operatorname{Sper}R\stackrel{{\scriptstyle\xi}}{{\rightarrow}}\operatorname{Spec}R,$
    where $\xi$ is defined by $\xi(\alpha):=\operatorname{Supp}(\alpha)$. (Note that if $R=F$ is a field, then $\operatorname{Sper}^{\mathrm{max}}R=\operatorname{Sper}R=X_{F}$, the space of orderings of $F$.)
The following theorems, due to Knebusch, explain how to obtain the maximal ordering associated to a signature on a semilocal ring. These results can be found in [13, Theorem 4.8] and [14, pp. 87-88]. The connection with maximal orderings is given in the second reference, but is presented for connected rings. It is pointed out that this assumption can be made without loss of generality, but we quickly present an argument in the proof below.
Theorem 1.18.
Assume that $R$ is semilocal and let $s$ be a signature on $R$. Define
$Q(s):=\Big{\{}r_{1}^{2}u_{1}+\cdots+r_{k}^{2}u_{k}\,\Big{|}\,k\in\mathbb{N},\ u_{i}\in R^{\times},\ s(\langle u_{i}\rangle)=1,\ \sum_{i=1}^{k}Rr_{i}=R\Big{\}}.$
Then
  1. (1)
    $Q(s)$ is closed under sum, $Q(s)\cap-Q(s)=\emptyset$ and $p(s):=R\setminus(Q(s)\cup-Q(s))$ is a prime ideal of $R$.
  2. (2)
    $\alpha(s):=Q(s)\cup p(s)$ is a maximal ordering on $R$ with support $p(s)$ such that $\operatorname{sign}_{\alpha(s)}=s$.
Proof.
For (1), see [13, Theorem 4.8]. For (2), it follows immediately from the properties of $Q(s)$ that $\alpha(s)$ is an ordering on $R$ with support $p(s)$ and that $\operatorname{sign}_{\alpha(s)}=s$. Suppose that $\alpha(s)$ is not maximal, so that there is $\beta\in\operatorname{Sper}R$ such that $\alpha(s)\subsetneq\beta$. Take $x\in\beta\setminus\alpha(s)$. Since $x\not\in\alpha(s)$ we have $x\in-\alpha(s)\setminus\alpha(s)=-Q(s)$ and thus $x=-(r_{1}^{2}u_{1}+\cdots+r_{k}^{2}u_{k})$ as described above. In $\mathrm{qf}(R/\operatorname{Supp}\beta)$ we have $\bar{x}=-\sum_{i=1}^{k}\bar{r}_{i}^{2}\bar{u}_{i}$ with $\bar{u}_{i}\in\bar{\beta}$, $\bar{u}_{i}\not=0$, so that $\bar{u}_{i}>_{\bar{\beta}}0$. In particular $\bar{x}\in-\bar{\beta}$. Since $\bar{x}\in\bar{\beta}$ by choice of $x$ we have $\bar{x}=0$, which implies that $\bar{r}_{i}=0$ for every $i$, i.e., $r_{1},\ldots,r_{k}\in\operatorname{Supp}\beta$, contradicting $\sum_{i=1}^{k}Rr_{i}=R$. ∎
Applying this result to $s=\operatorname{sign}_{\alpha}$ for $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$, we obtain
Theorem 1.19.
Assume that $R$ is semilocal and that $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$. Then
$\alpha\setminus\operatorname{Supp}(\alpha)=\Big{\{}r_{1}^{2}u_{1}+\cdots+r_{k}^{2}u_{k}\,\Big{|}\,k\in\mathbb{N},\ u_{i}\in\alpha\cap R^{\times},\ \sum_{i=1}^{k}Rr_{i}=R\Big{\}}.$
The special case of $R$ local easily follows, but is worth noting and was obtained earlier (without the link to $\operatorname{Sper}R$, which was introduced later), first in [12] if $2\in R^{\times}$, and then in general in [13]:
$\alpha\setminus\operatorname{Supp}(\alpha)=\{u_{1}+\cdots+u_{k}\mid k\in\mathbb{N},\ u_{i}\in\alpha\cap R^{\times}\}.$
Finally, we mention the following simple fact for future use:
Lemma 1.20.
Let $F$ be a field with space of orderings $X_{F}$. Let $P\in X_{F}$, and denote by $F_{P}$ any real closure of $F$ at $P$. Then $F$ is cofinal in $F_{P}$, i.e., for every $a\in F_{P}$ there exists $b\in F$ such that $a\leq b$. In particular, for every $a\in F_{P}$, if $a>0$, then there exists $b\in F$ such that $0<b\leq a$.
Proof.
Let $p(X)=a_{0}+a_{1}X+\cdots+a_{k-1}X^{k-1}+X^{k}\in F[X]$ be such that $p(a)=0$. Then $a\leq\max\{1,|a_{0}|+\cdots+|a_{k-1}|\}$ by [15, Proposition 1.7.1]. The second statement follows by taking inverses. ∎
1.5. Positive definite matrices
The results in this section are well-known, but we could not find a reference for the quaternion case. In this section we assume that $F$ is a real closed field and that $(D,\vartheta)\in\{(F,\operatorname{id}),(F(\sqrt{-1}),\gamma),((-1,-1)_{F},\gamma)\}$, where $\gamma$ denotes the canonical involution in each case. We consider norms with values in $F$. For vectors in $D^{k}$ we consider the euclidean norm, i.e., for $X=(x_{1},\ldots,x_{k})^{t}\in D^{k}$ we have $\|X\|=\sqrt{\sum_{i=1}^{k}n(x_{i})}$, where $n(x):=\vartheta(x)x$. On $M_{k}(D)$ we use the induced operator norm, i.e., $\|M\|_{\mathrm{op}}:=\sup_{\|X\|=1}\|MX\|$. Tarski’s transfer principle [24, Corollary 11.5.4] ensures that this supremum exists and that the operator norm is equivalent to the maximum norm determined by the unique ordering of $F$ on the $F$-vector space $M_{k}(D)$ (both properties can be expressed by first-order formulas in the language of ordered fields, that are true in $\mathbb{R}$). Therefore, $\|\cdot\|_{\mathrm{op}}$ defines the same topology as the one induced by the ordering of $F$.
Lemma 1.21.
Let $k\in\mathbb{N}$. Then
$\operatorname{PD}_{k}(D,\vartheta):=\{B\in\operatorname{Sym}(M_{k}(D),\vartheta^{t})\mid\vartheta(X)^{t}BX>0\text{ for every }X\in D^{k}\setminus\{0\}\}$
is an open subset of $\operatorname{Sym}(M_{k}(D),\vartheta^{t})$ for the topology induced by the unique ordering of $F$.
Proof.
Since this property can be expressed by a first-order formula in the language of fields in each of the three cases $(D,\vartheta)\in\{(F,\operatorname{id}),(F(\sqrt{-1}),\gamma),((-1,-1)_{F},\gamma)\}$, it suffices to prove it for $F=\mathbb{R}$ by Tarski’s transfer principle [24, Corollary 11.5.4]. The well-known proof below works in all three cases, and we give the details in order to point out that it also works in the quaternion case. Observe that for every $X,Y\in D^{k}$, we have $\|\vartheta(X)^{t}Y\|\leq\|X\|\cdot\|Y\|$ (the verification is direct in the quaternion case).
We reformulate $M\in\operatorname{PD}_{k}(D,\vartheta)$ as: there is $\delta>0$ such that $\vartheta(X)^{t}MX\geq\delta$ for every $X\in D^{k}$, $\|X\|=1$. Let $M\in\operatorname{PD}_{k}(D,\vartheta)$ and let $\delta$ be as described above. Let $B\in\operatorname{Sym}(M_{k}(D),\vartheta^{t})$ be such that $\|B-M\|_{\mathrm{op}}\leq\varepsilon$. Then
$\vartheta(X)^{t}BX=\vartheta(X)^{t}MX+(\vartheta(X)^{t}BX-\vartheta(X)^{t}MX),$
and, if $\|X\|=1$,
$|\vartheta(X)^{t}BX-\vartheta(X)^{t}MX|$ $=|\vartheta(X)^{t}(B-M)X|=\|\vartheta(X)^{t}(B-M)X\|$
$\leq\|X\|\cdot\|(B-M)X\|\leq\|B-M\|_{\mathrm{op}}\leq\varepsilon$
from which the result follows if we take $\varepsilon=\delta/2$. ∎
2. $\mathscr{M}$-Signatures of hermitian forms
If $(A,\sigma)$ is a central simple $F$-algebra with involution over a field $F$, $h$ is a hermitian form over $(A,\sigma)$, and $P$ is an ordering on $F$, we defined in [1, Section 3.2] the signature of $h$ at $P$ with respect to a particular Morita equivalence $\mathscr{M}_{P}$ (which determines the sign of the signature), denoted $\operatorname{sign}^{\mathscr{M}_{P}}_{P}h$. We only considered nonsingular forms in [1, Section 3.2], which was unnecessarily restrictive since the method we used (reduction to the Sylvester signature via scalar extension to a real closure of $F$ at $P$ and hermitian Morita theory) applies in fact to forms that may be singular.
We will use the notation and results from [1]. Note that this signature is also presented in [2, second half of p. 499], where the omission in [1] of one (irrelevant) case for $(D_{P},\vartheta_{P})$ has been rectified.
Let $R$ be a commutative ring (with $2\in R^{\times}$), let $(A,\sigma)$ be an Azumaya algebra with involution over $R$, let $S=Z(A)$ and let $\iota:=\sigma|_{S}$.
Definition 2.1.
Let $h$ be a hermitian form over $(A,\sigma)$ and let $\alpha\in\operatorname{Sper}R$. Then $h\otimes\kappa(\alpha)$ is a hermitian form over the central simple algebra with involution $(A(\alpha),\sigma(\alpha))$ and we define the $\mathscr{M}$-signature of $h$ at $\alpha$ by
(2.1) $\operatorname{sign}^{\mathscr{M}}_{\alpha}h:=\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(h\otimes\kappa(\alpha)),$
where $\mathscr{M}_{\bar{\alpha}}$ is a Morita equivalence as in [1, Section 3.2].
Note that the superscript $\mathscr{M}$ on the left hand side of (2.1) signifies that each computation of $\operatorname{sign}^{\mathscr{M}}_{\alpha}$ depends on a choice of Morita equivalence $\mathscr{M}_{\bar{\alpha}}$. The use of a different Morita equivalence can result at most in a change of sign, cf. [1, Proposition 3.4]. Therefore, the notation $\operatorname{sign}^{\mathscr{M}}_{\alpha}$ should specify what Morita equivalence $\mathscr{M}_{\bar{\alpha}}$ is used but, in order not to overload the notation, we assume that for a given $\alpha$ we always use the same Morita equivalence $\mathscr{M}_{\bar{\alpha}}$.
In fact, the main drawback of the $\mathscr{M}$-signature is that the sign of the signature of a form can be changed arbitrarily at each ordering by taking a different Morita equivalence. This is in particular a problem if we hope to consider the total signature of a form as a continuous function on $\operatorname{Sper}R$. We solved this problem in the case of central simple algebras with involution by introducing a “reference form”, that determines the sign of the signature at each ordering, cf. [1, Section 6] and [2, Section 3]. We will show in a forthcoming publication that the same can be done in the case of Azumaya algebras with involution.
Remark 2.2.
Let $(B,\tau)$ be a central simple $F$-algebra with involution, $P\in X_{F}$, and $F_{P}$ a real closure of $F$ at $P$. Note that $F=\operatorname{Sym}(Z(B),\tau)$. Then $B\cong M_{n}(D)$ and $B\otimes_{F}F_{P}\cong M_{n_{P}}(D_{P})$, where $D$ and $D_{P}$ are division algebras with involution over $F$ and $F_{P}$, respectively. (We do not need to name the involutions on $D$ and $D_{P}$ here.) Clearly, $n\leq n_{P}\leq\sqrt{\dim_{F}B}$. Let $b\in\operatorname{Sym}(B,\tau)$. By [3, Proposition 4.4], we have $\operatorname{sign}^{\mathscr{M}_{P}}_{P}\langle b\rangle_{\tau}\leq n_{P}\leq\sqrt{\dim_{F}B}$. Applying this to $(A(\alpha),\sigma(\alpha))$ in Definition 2.1, let $\mathcal{S}$ be a set of generators of $A$ as an $R$-module, and let $a\in\operatorname{Sym}(A,\sigma)$. We have
$\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}\leq\sqrt{\dim_{\kappa(\alpha)}A(\alpha)}\leq|\mathcal{S}|,$
which provides a bound on $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}$ which is independent of $a$ and $\alpha$.
Proposition 2.3.
Let $h$ be a hermitian form over $(A,\sigma)$, let $q$ be a quadratic form over $R$ and let $\alpha\in\operatorname{Sper}R$. Then
$\operatorname{sign}^{\mathscr{M}}_{\alpha}(q\otimes h)=(\operatorname{sign}_{\alpha}q)\cdot(\operatorname{sign}^{\mathscr{M}}_{\alpha}h),$
where the same Morita equivalence is used in the computation of the signature on both sides of the equality.
Proof.
By definition $\operatorname{sign}^{\mathscr{M}}_{\alpha}(q\otimes h)$ is equal to $\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}((q\otimes h)\otimes\kappa(\alpha))$, where $(q\otimes h)\otimes\kappa(\alpha)$ is considered as a form over $(A(\alpha),\sigma(\alpha))$. But $\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}((q\otimes h)\otimes\kappa(\alpha))=(\operatorname{sign}_{\bar{\alpha}}q\otimes\kappa(\alpha))\cdot(\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}h\otimes\kappa(\alpha))$ by [1, Proposition 3.6]. The result follows. ∎
Lemma 2.4.
Let $T$ be a quadratic étale $R$-algebra with standard involution $\vartheta$. Let $h$ be a hermitian form over $(T,\vartheta)$. Then, for every $\alpha\in\operatorname{Sper}R$, $\operatorname{sign}^{\mathscr{M}}_{\alpha}h=0$ implies $\operatorname{sign}_{\alpha}\operatorname{Tr}_{T/R}(h)=0$.
Proof.
Note that $(T,\vartheta)$ is an Azumaya algebra with involution over $R$ and that $\vartheta\otimes\operatorname{id}$ is the standard involution on $T\otimes_{R}k(\alpha)$ by Remark 1.9. In particular, signatures are defined, and by the definition of signatures for central simple algebras with involution, the form $h\otimes_{R}k(\alpha)$ has signature 00 at the unique ordering on $k(\alpha)$. Writing $h\otimes_{R}k(\alpha)\simeq\varphi\perp 0$, where $\varphi$ is nonsingular and 00 is the zero form of appropriate rank, cf. [3, Proposition A.3], it follows that $\varphi$ has signature 00 at the unique ordering on $k(\alpha)$, and thus is weakly hyperbolic (i.e., ×$\ell\times\varphi$ is hyperbolic for some $\ell\in\mathbb{N}$) by [21, Theorem 4.1] (or [7, Theorem 6.5]).
As recalled in Section 1.2, $\operatorname{Tr}_{T/R}$ is $R$-linear and $\operatorname{Tr}_{T/R}(h)$ is a quadratic form over $R$. Also, $\operatorname{Tr}_{T\otimes_{R}k(\alpha)/k(\alpha)}$ is an involution trace for $\vartheta\otimes\operatorname{id}$, and it follows that $\operatorname{Tr}_{T\otimes_{R}k(\alpha)/k(\alpha)}(\varphi)$ is weakly hyperbolic by Section 1.2, and thus has signature 00. Since $\operatorname{Tr}_{T/R}(h)\otimes k(\alpha)=\operatorname{Tr}_{T\otimes_{R}k(\alpha)/k(\alpha)}(h\otimes k(\alpha))$ by Remark 1.9, and observing that
$\operatorname{Tr}_{T\otimes_{R}k(\alpha)/k(\alpha)}(h\otimes k(\alpha))\simeq\operatorname{Tr}_{T\otimes_{R}k(\alpha)/k(\alpha)}(\varphi)\perp\operatorname{Tr}_{T\otimes_{R}k(\alpha)/k(\alpha)}(0),$
it follows that $\operatorname{sign}_{\alpha}\operatorname{Tr}_{T/R}(h)=0$. ∎
Definition 2.5.
We call
$\operatorname{Nil}[A,\sigma]:=\{\alpha\in\operatorname{Sper}R\mid\operatorname{sign}^{\mathscr{M}}_{\alpha}=0\}$
the set of nil orderings of $(A,\sigma)$. By the observation after Definition 2.1, $\operatorname{Nil}[A,\sigma]$ is independent of the choice of Morita equivalence at each $\alpha$.
Remark 2.6.
If $(A,\sigma)$ is a central simple $F$-algebra with involution, this definition is equivalent to our original one ([1, Definition 3.7] by [1, Theorem 6.4]), from which it follows that if $\sigma$ is orthogonal and $\tau$ is any symplectic involution on $A$, then $\operatorname{Nil}[A,\sigma]=X_{F}\setminus\operatorname{Nil}[A,\tau]$.
Also note that if $P\in X_{F}\setminus\operatorname{Nil}[A,\sigma]$ and $F_{P}$ is a real closure of $F$ at $P$, then $(A\otimes_{F}F_{P},\sigma\otimes\operatorname{id})\cong(M_{n_{P}}(D_{P}),\operatorname{Int}(\Phi_{P})\circ{\vartheta_{P}}^{t})$, where $D_{P}\in\{F_{P},F_{P}(\sqrt{-1}),(-1,-1)_{F_{P}}\}$, $\vartheta_{P}$ is the canonical involution on $D_{P}$, and $\Phi_{P}\in\operatorname{Sym}(M_{n_{P}}(D_{P})^{\times},{\vartheta_{P}}^{t})$, cf. [4, p. 4 and Remark 6.2].
Lemma 2.7.
Let $\alpha\in\operatorname{Sper}R$. Then statements (1) and (2) below are equivalent:
  1. (1)
    $\alpha\in\operatorname{Nil}[A,\sigma]$;
  2. (2)
    $\bar{\alpha}\in\operatorname{Nil}[A(\alpha),\sigma(\alpha)]$.
Assume in addition that $\sigma$ is of unitary type at $\operatorname{Supp}(\alpha)$. Then (1) and (2) and the following statements are equivalent:
  1. (3)
    $Z(A(\alpha))\otimes_{\kappa(\alpha)}k(\bar{\alpha})\cong k(\bar{\alpha})\times k(\bar{\alpha})$, i.e., $Z(A(\alpha))\otimes_{\kappa(\alpha)}k(\alpha)\cong k(\alpha)\times k(\alpha)$ since we can take $k(\bar{\alpha})=k(\alpha)$ as observed before;
  2. (4)
    $Z(A)\otimes_{R}k(\alpha)\cong k(\alpha)\times k(\alpha)$;
  3. (5)
    $\alpha\in\operatorname{Nil}[S,\iota]$.
Proof.
(1$\Rightarrow$(2): Assume $\bar{\alpha}\not\in\operatorname{Nil}[A(\alpha),\sigma(\alpha)]$. Then there exists $z\in\operatorname{Sym}(A(\alpha),\sigma(\alpha))$ such that $\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle z\rangle_{\sigma(\alpha)}\not=0$ (see [1, Theorem 6.4]). Using Remark 1.17, write $z=a\otimes\frac{1}{\bar{r}}$, for some $r\in R\setminus\operatorname{Supp}(\alpha)$. Since $\bar{r}$ is invertible in $\kappa(\alpha)$, we have
$0\not=\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle z\rangle_{\sigma(\alpha)}=\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle z{\bar{r}}^{2}\rangle_{\sigma(\alpha)}=\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle ar\otimes 1\rangle_{\sigma\otimes\operatorname{id}_{\kappa(\alpha)}}=\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle ar\rangle_{\sigma},$
contradicting that $\alpha\in\operatorname{Nil}[A,\sigma]$.
(2$\Rightarrow$(1): This follows from the fact that $\operatorname{sign}^{\mathscr{M}}_{\alpha}h=\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}h\otimes\kappa(\alpha)=0$ for any hermitian form $h$ over $(A,\sigma)$.
For the remaining equivalences, recall that if $(B,\tau)$ is a central simple algebra with involution of the second kind over a field $F$, then $P\in\operatorname{Nil}[B,\tau]$ if and only if $Z(B\otimes F_{P})\cong F_{P}\times F_{P}$, cf. [2, bottom of p. 499 and Def. 2.1].
The equivalence (2$\Leftrightarrow$(3) follows from this observation, and we have (3$\Leftrightarrow$(4) since $k(\alpha)$ is a real closure of $\kappa(\alpha)$ at $\bar{\alpha}$ and $Z(A(\alpha))=Z(A)\otimes_{R}\kappa(\alpha)$ (by Proposition 1.7). Finally, the equivalence (4$\Leftrightarrow$(5) follows from (1$\Leftrightarrow$(4) applied to $(S,\iota)$ since $S=Z(S)=Z(A)$. ∎
Corollary 2.8.
Assume that $R$ is semilocal and that $\sigma$ is of unitary type. Then there is $d\in R^{\times}$ such that $\operatorname{Nil}[A,\sigma]=\mathring{H}(d)$.
Proof.
By [8, Proposition 1.21], $S=Z(A)$ is a quadratic étale $R$-algebra, and for every $\alpha\in\operatorname{Sper}R$, $\sigma(\alpha)$ is of unitary type at the prime ideal $\operatorname{Supp}(\alpha)$. Furthermore, as recalled in Section 1.2, $S=R\oplus\lambda R$ for some $\lambda\in S$ such that $d:=\lambda^{2}\in R^{\times}$. In particular, for any $\alpha\in\operatorname{Sper}R$, $Z(A(\alpha))=S\otimes_{R}\kappa(\alpha)=\kappa(\alpha)\oplus(\lambda\otimes 1)\kappa(\alpha)=\kappa(\alpha)(\sqrt{\bar{d}})$, where $\bar{d}$ is the image of $d$ in $R/\operatorname{Supp}(\alpha)$, and where the first equality follows from Proposition 1.7. Therefore, for $\alpha\in\operatorname{Sper}R$:
[…]
$\Leftrightarrow\kappa(\alpha)(\sqrt{\bar{d}})\otimes_{\kappa(\alpha)}k(\bar{\alpha})\cong k(\bar{\alpha})\times k(\bar{\alpha})$
$\Leftrightarrow\sqrt{\bar{d}}\in k(\bar{\alpha})\Leftrightarrow\bar{d}\in\bar{\alpha}\Leftrightarrow d\in\alpha$
$\Leftrightarrow\alpha\in\mathring{H}(d)\ \text{since }d\text{ is invertible}.\qed$
Remark 2.9.
Recall from Section 1.2 that when $R$ is connected, then either $S$ is connected or $S\cong R\times R$. Moreover, if $\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]\not=\emptyset$, we cannot have $S\cong R\times R$ by Lemma 2.7. Therefore, if $R$ is connected and $\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]\not=\emptyset$, then $S$ is also connected.
Lemma 2.10.
Let $F$ be a field and let $(B,\tau)$ be a central simple $F$-algebra with involution of the first kind. If $\operatorname{Nil}[B,\tau]\not=\emptyset$, then $\deg B$ is even.
Proof.
Assume that $\deg B$ is odd. Then $B$ is split and $\tau$ is orthogonal by [17, Corollary 2.8(1)]. It follows that $\operatorname{Nil}[B,\tau]=\emptyset$ by [1, Definition 3.7]. ∎
Lemma 2.11.
Assume that $R$ is semilocal connected and that $\sigma$ is of orthogonal or symplectic type. If $\operatorname{Nil}[A,\sigma]\not=\emptyset$, then $\operatorname{Skew}(A^{\times},\sigma)\not=\emptyset$.
Proof.
Let $\alpha\in\operatorname{Nil}[A,\sigma]$. Then $\bar{\alpha}\in\operatorname{Nil}[A(\alpha),\sigma(\alpha)]$ by Lemma 2.7. Hence, $\deg A\otimes_{R}\kappa(\alpha)$ is even by Lemma 2.10. Since $R$ is connected, the rank of $A$ is constant (since $A$ is a projective $R$-module, see [26, p. 12]). It follows that $\deg A\otimes_{R}\mathrm{qf}(R/\mathfrak{p})$ is even for every $\mathfrak{p}\in\operatorname{Spec}R$. We can then apply [8, Lemma 1.26] with $\varepsilon=-1$. ∎
2.1. Elements of maximal signature
In [4], working with central simple algebras with involution, we investigated the maximal value that the signature at $P\in X_{F}$ can take (when it is non-zero) when applied to one-dimensional nonsingular forms. We found that this maximal value is the matrix size of the algebra over its skew-field part after scalar extension to $F_{P}$ (and linked it to the existence of positive involutions), cf. [4, Proposition 6.7, Theorem 6.8].
We are interested in the same question when $(A,\sigma)$ is an Azumaya algebra with involution. More precisely, we will show in Corollary 2.19 that if $\alpha\not\in\operatorname{Nil}[A,\sigma]$, then this maximal value is the matrix size $n_{\bar{\alpha}}$ of $A\otimes_{R}k(\alpha)$ over its skew-field part (i.e., $A\otimes_{R}k(\alpha)\cong M_{n_{\bar{\alpha}}}(D_{\bar{\alpha}})$ using the notation from Remark 2.2). We first introduce some notation:
Definition 2.12.
If $(B,\tau)$ is an Azumaya algebra with involution over $R$, and $\alpha\in\operatorname{Sper}R$, we define:
$m_{\alpha}(B,\tau)$ $:=\max\{\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle b\rangle_{\tau}\mid b\in\operatorname{Sym}(B^{\times},\tau)\}$
and
$M^{\mathscr{M}}_{\alpha}(B,\tau)$ :=$:=\{b\in\operatorname{Sym}(B^{\times},\tau)\mid\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle b\rangle_{\tau}=m_{\alpha}(B,\tau)\}.$
Observe that $m_{\alpha}(B,\tau)$ is independent of the choice of the Morita equivalence $\mathscr{M}_{\bar{\alpha}}$ (cf. Definition 2.1), and is finite by Remark 2.2.
We introduce some notation that will be used in the next four results. For $\alpha\in\operatorname{Sper}R$, define
$\mathscr{S}_{\alpha}(A,\sigma):=\bigcup\{D_{(A,\sigma)}\langle a_{1},\ldots,a_{k}\rangle_{\sigma}\mid k\in\mathbb{N},\ a_{i}\in$ $\operatorname{Sym}(A,\sigma),$
$a_{i}\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))\}.$
Furthermore, for $\mathfrak{p}\in\operatorname{Spec}R$, we denote by $\pi_{\mathfrak{p}}$ the canonical projection from $R$ to $R/\mathfrak{p}$ and by $\pi_{A\mathfrak{p}}$ the canonical projection from $A$ to $A/A\mathfrak{p}$. Then, denoting by $\bar{\sigma}$ the involution induced by $\sigma$ on $A/A\mathfrak{p}$, we define
$\mathscr{S}^{\prime}_{\mathfrak{p}}:=\bigcup\{D_{(A/A\mathfrak{p},\bar{\sigma})}\langle\pi_{A\mathfrak{p}}(a_{1}),\ldots,\pi_{A\mathfrak{p}}(a_{k})\rangle_{\bar{\sigma}}\mid k\in\mathbb{N},$ $a_{i}\in\operatorname{Sym}(A,\sigma),$
$a_{i}\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))\}.$
Lemma 2.13.
Let $\alpha\in\operatorname{Sper}R$. Then there is an element $b\in\operatorname{Sym}(A,\sigma)$ such that $b\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. In particular $b\in\mathscr{S}_{\alpha}(A,\sigma)$.
Proof.
Let $c\in\operatorname{Sym}(A(\alpha)^{\times},\sigma(\alpha))$ be such that $\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle c\rangle_{\sigma(\alpha)}=m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. By Remark 1.17, there are $b_{0}\in A$ and $r\in R\setminus\operatorname{Supp}(\alpha)$ such that $c=b_{0}\otimes\frac{1}{\bar{r}}$. Let $b_{1}:=\frac{1}{2}(b_{0}+\sigma(b_{0}))$. Then $b_{1}\in\operatorname{Sym}(A,\sigma)$ and
$b_{1}\otimes\frac{1}{\bar{r}}=\frac{1}{2}(b_{0}\otimes\frac{1}{\bar{r}}+\sigma(b_{0})\otimes\frac{1}{\bar{r}})=\frac{1}{2}(c+\sigma(\alpha)(c))=c.$
We take $b:=rb_{1}$. Then $b\in\operatorname{Sym}(A,\sigma)$ and, since $\bar{r}\in\kappa(\alpha)^{\times}$, we have $b\otimes 1=\bar{r}^{2}c\in A(\alpha)^{\times}$, and
$\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle b\otimes 1\rangle_{\sigma(\alpha)}=\operatorname{sign}^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}\langle\bar{r}^{2}c\rangle_{\sigma(\alpha)}=m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha)),$
so that $b\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. ∎
Lemma 2.14.
Assume that $R=F$ is a field and let $P\in X_{F}\setminus\operatorname{Nil}[A,\sigma]$. Let $a\in\operatorname{Sym}(A,\sigma)$ be such that $\operatorname{sign}^{\mathscr{M}}_{P}\langle a\rangle_{\sigma}=m_{P}(A,\sigma)$. Then:
  1. (1)
    $a$ is invertible in $A$;
  2. (2)
    There is $\mu\in P\setminus\{0\}$ such that $\operatorname{sign}^{\mathscr{M}}_{P}\langle a-r\rangle_{\sigma}=m_{P}(A,\sigma)$ for every $r\in P$ such that $r\leq_{P}\mu$.
Proof.
By [4, Proposition 6.7] we have $m_{P}(A,\sigma)=n_{P}$. We use the notation from Remark 2.6. Since $P\not\in\operatorname{Nil}[A,\sigma]$, it follows from the computation of $\mathscr{M}$-signatures (cf. the beginning of Section 2) that $\operatorname{sign}^{\mathscr{M}}_{P}\langle a\rangle_{\sigma}$ is equal to the Sylvester signature of the form $\langle\Phi_{P}^{-1}(a\otimes 1)\rangle_{{\vartheta_{P}}^{t}}$, where $\Phi_{P}^{-1}(a\otimes 1)\in\operatorname{Sym}(M_{n_{P}}(D_{P}),{\vartheta_{P}}^{t})$. Since $D_{P}\in\{F_{P},F_{P}(\sqrt{-1}),(-1,-1)_{F_{P}}\}$ and $\vartheta_{P}$ is the canonical involution on $D_{P}$, the matrix $\Phi_{P}^{-1}(a\otimes 1)$ can be diagonalized by congruences (which does not change the Sylvester signature), so we can assume that $\Phi_{P}^{-1}(a\otimes 1)$ is diagonal. Since it has Sylvester signature $n_{P}$, its diagonal elements are all positive, i.e., $\Phi_{P}^{-1}(a\otimes 1)\in\operatorname{PD}_{n_{P}}(D_{P},\vartheta_{P})$. In particular, $\Phi_{P}^{-1}(a\otimes 1)$ is invertible. Therefore $a\otimes 1$ is not a zero divisor, and $a$ is not a zero divisor in $A$. It follows that $a$ is invertible since $A$ is Artinian. This proves (1).
For (2): The element $a\otimes 1$ is in $\Phi_{P}\cdot\operatorname{PD}_{n_{P}}(D_{P},\vartheta_{P})$, which is an open subset of $\operatorname{Sym}(M_{n_{P}}(D_{P}),{\vartheta_{P}}^{t})$ by Lemma 1.21. It follows that there is $\varepsilon>0$ in $F_{P}$ such that for all $M\in\operatorname{Sym}(M_{n_{P}}(D_{P}),{\vartheta_{P}}^{t})$ that satisfy $\|M\|_{\mathrm{op}}<\varepsilon$ we have $a\otimes 1-M\in\Phi_{P}\cdot\operatorname{PD}_{n_{P}}(D_{P},\vartheta_{P})$. Taking $\mu\in F_{P}$ such that $0<\mu<\varepsilon$, we obtain $\|\mu I_{n_{P}}\|_{\mathrm{op}}<\varepsilon$. Hence, $a\otimes 1-\mu I_{n_{P}}\in\Phi_{P}\cdot\operatorname{PD}_{n_{P}}(D_{P},\vartheta_{P})$. In particular, the signature of the form $\langle a\otimes 1-\mu I_{n_{P}}\rangle_{\sigma\otimes\operatorname{id}}$ equals $n_{P}$.
However, $F$ is cofinal in $F_{P}$ by Lemma 1.20, so we can find such a $\mu$ in $F$. The choice of $\mu$ guarantees that $\operatorname{sign}^{\mathscr{M}}_{P}\langle a-r\rangle_{\sigma}=n_{P}=m_{P}(A,\sigma)$ whenever $r\in P$, $r\leq_{P}\mu$. ∎
Lemma 2.15.
Let $\mathfrak{m}$ be a maximal ideal of $R$ and let $\alpha\in\operatorname{Sper}R$ be such that $\operatorname{Supp}\alpha\subseteq\mathfrak{m}$. Assume that
(2.2) $\forall a\in\operatorname{Sym}(A,\sigma)\quad a\otimes_{R}1_{\kappa(\alpha)}\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))\text{ implies }a\in A\mathfrak{m}.$
Then property (2.2) is preserved under quotients by $\operatorname{Supp}\alpha$ and, if $\operatorname{Supp}\alpha=\{0\}$, under localization at $\mathfrak{m}$. More precisely:
  1. (1)
    Let $R_{1}:=R/\operatorname{Supp}\alpha$. Then property (2.2) holds for $(A_{1},\sigma_{1}):=(A\otimes_{R}R_{1},\sigma\otimes_{R}\operatorname{id}_{R_{1}})$ and the ordering $\alpha_{1}$ induced by $\alpha$ on $R_{1}$ together with the maximal ideal $\mathfrak{m}_{1}:=\mathfrak{m}/\operatorname{Supp}\alpha$ of $R_{1}$, i.e.,
    $\forall a\in\operatorname{Sym}(A_{1},\sigma_{1})\quad a\otimes_{R_{1}}1_{\kappa(\alpha_{1})}\in M^{\mathscr{M}_{\bar{\alpha}_{1}}}_{\bar{\alpha}_{1}}(A_{1}(\alpha_{1}),\sigma_{1}(\alpha_{1}))\text{ implies }a\in A_{1}\mathfrak{m}_{1}.$
    Furthermore, if $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$, then $\alpha_{1}\in\operatorname{Sper}^{\mathrm{max}}R_{1}$.
  2. (2)
    Assume that $\operatorname{Supp}\alpha=\{0\}$. Then property (2.2) holds for $(A_{2},\sigma_{2}):=(A\otimes_{R}R_{\mathfrak{m}},\sigma\otimes_{R}\operatorname{id}_{R_{\mathfrak{m}}})$ and the ordering $\alpha_{2}$ induced by $\alpha$ on $R_{\mathfrak{m}}$ together with the unique maximal ideal $\mathfrak{m}_{2}$ of $R_{\mathfrak{m}}$, i.e.,
    $\forall a\in\operatorname{Sym}(A_{2},\sigma_{2})\quad a\otimes_{R_{\mathfrak{m}}}1_{\kappa(\alpha_{2})}\in M^{\mathscr{M}_{\bar{\alpha}_{2}}}_{\bar{\alpha}_{2}}(A_{2}(\alpha_{2}),\sigma_{2}(\alpha_{2}))\text{ implies }a\in A_{2}\mathfrak{m}_{2}.$
    Furthermore, if $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$, then $\alpha_{2}\in\operatorname{Sper}^{\mathrm{max}}R_{\mathfrak{m}}$.
Proof.
(1) We have natural maps
$R\rightarrow R_{1}=R/\operatorname{Supp}\alpha\rightarrow\mathrm{qf}(R_{1})=\kappa(\alpha)=\kappa(\alpha_{1})$
with $\bar{\alpha}=\bar{\alpha}_{1}$, and thus
$A\rightarrow A_{1}=A\otimes_{R}R_{1}\rightarrow A\otimes_{R}\mathrm{qf}(R_{1})=A(\alpha)=A_{1}(\alpha_{1}),$
while
$\sigma_{1}(\alpha_{1}):=\sigma_{1}\otimes_{R_{1}}\operatorname{id}_{\kappa(\alpha_{1})}=(\sigma\otimes_{R}\operatorname{id}_{R_{1}})\otimes_{R_{1}}\operatorname{id}_{\kappa(\alpha_{1})}=\sigma\otimes_{R}\operatorname{id}_{\kappa(\alpha_{1})}=\sigma(\alpha).$
Let $b\in\operatorname{Sym}(A_{1},\sigma_{1})$ be such that
$b\otimes_{R_{1}}1_{\kappa(\alpha_{1})}\in M^{\mathscr{M}_{\bar{\alpha}_{1}}}_{\bar{\alpha}_{1}}(A_{1}(\alpha_{1}),\sigma_{1}(\alpha_{1}))=M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha)).$
Then $b=c\otimes_{R}1_{R_{1}}$ for some $c\in A$ (the argument is similar to Remark 1.17) and
$b\otimes_{R_{1}}1_{\kappa(\alpha_{1})}=c\otimes_{R}1_{R_{1}}\otimes_{R_{1}}1_{\kappa(\alpha_{1})}=c\otimes_{R}1_{\kappa(\alpha)}.$
Therefore $c\otimes_{R}1_{\kappa(\alpha)}\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$ and thus $c\in A\mathfrak{m}$ by (2.2). It follows that $b=c\otimes_{R}1_{R_{1}}\in A\mathfrak{m}\otimes_{R}1_{R_{1}}\subseteq(A\otimes_{R}R_{1})(\mathfrak{m}/\operatorname{Supp}\alpha)=A_{1}\mathfrak{m}_{1}$ (the inclusion follows from $am\otimes 1=(a\otimes 1)(m\otimes 1)=(a\otimes 1)(1\otimes(m+\operatorname{Supp}\alpha))$).
The statement about the maximality of $\alpha_{1}$ follows from the fact that the homeomorphism in [15, Proposition 3.3.11] clearly preserves inclusions.
(2) Note that $\alpha_{2}$ is indeed an ordering on $R_{\mathfrak{m}}$ since $(R\setminus\mathfrak{m})\cap\operatorname{Supp}\alpha=\emptyset$, cf. [15, Proposition 3.3.10]. We also have
$\alpha_{2}=\Big{\{}\dfrac{r}{s^{2}}\,\Big{|}\,r\in\alpha,\ s\in R\setminus\mathfrak{m}\Big{\}}$
(by [15, Proof of Proposition 3.3.10]) and $\operatorname{Supp}\alpha_{2}=\{0\}$, so that $\kappa(\alpha_{2})=\mathrm{qf}(R_{\mathfrak{m}})$. Observe that the map $R\rightarrow\mathrm{qf}(R)$, which is the first step in the computation of signatures of elements of $\operatorname{Sym}(A,\sigma)$ (since $\operatorname{Supp}\alpha=\{0\}$, cf. Definition 2.1) factors through $R_{\mathfrak{m}}$, giving $R\rightarrow R_{\mathfrak{m}}\rightarrow\mathrm{qf}(R)$, with $\mathrm{qf}(R)=\mathrm{qf}(R_{\mathfrak{m}})$, i.e., $\kappa(\alpha)=\kappa(\alpha_{2})$. Finally, a direct verification shows that $\bar{\alpha}=\bar{\alpha}_{2}$.
Let $b\in\operatorname{Sym}(A_{2},\sigma_{2})$ be such that $b\otimes_{R_{\mathfrak{m}}}1_{\kappa(\alpha_{2})}\in M^{\mathscr{M}_{\bar{\alpha}_{2}}}_{\bar{\alpha}_{2}}(A_{2}(\alpha_{2}),\sigma_{2}(\alpha_{2}))$. Then $b=c\otimes_{R}\dfrac{1}{s}$ for some $s\in R\setminus\mathfrak{m}$ (the argument is again similar to Remark 1.17). Since $s$ is invertible in $R_{\mathfrak{m}}$, $b$ has the same signature at $\alpha_{2}$ as $bs^{2}=cs\otimes_{R}1_{R_{\mathfrak{m}}}$, so that $(cs\otimes_{R}1_{R_{\mathfrak{m}}})\otimes_{R_{\mathfrak{m}}}1_{\kappa(\alpha_{2})}\in M^{\mathscr{M}_{\bar{\alpha}_{2}}}_{\bar{\alpha}_{2}}(A_{2}(\alpha_{2}),\sigma_{2}(\alpha_{2}))$.
We have $(cs\otimes_{R}1_{R_{\mathfrak{m}}})\otimes_{R_{\mathfrak{m}}}1_{\kappa(\alpha_{2})}=cs\otimes_{R}1_{\kappa(\alpha)}$, and thus
$cs\otimes_{R}1_{\kappa(\alpha)}\in$ $M^{\mathscr{M}_{\bar{\alpha}_{2}}}_{\bar{\alpha}_{2}}(A_{2}(\alpha_{2}),\sigma_{2}(\alpha_{2}))$
$=M^{\mathscr{M}_{\bar{\alpha}_{2}}}_{\bar{\alpha}_{2}}((A\otimes_{R}R_{\mathfrak{m}})\otimes_{R_{\mathfrak{m}}}\mathrm{qf}(R_{\mathfrak{m}}),(\sigma\otimes\operatorname{id}_{R_{\mathfrak{m}}})\otimes\operatorname{id}_{\kappa(\alpha_{2})})$
$=M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A\otimes_{R}\kappa(\alpha),\sigma\otimes\operatorname{id}_{\kappa(\alpha)}).$
By property (2.2), we obtain $cs\in A\mathfrak{m}$. Therefore $cs\otimes_{R}1_{R_{\mathfrak{m}}}\in A\mathfrak{m}\otimes_{R}1_{R_{\mathfrak{m}}}\subseteq(A\otimes_{R}R_{\mathfrak{m}})\mathfrak{m}_{2}$ (the inclusion follows from $am\otimes_{R}1_{R_{\mathfrak{m}}}=(a\otimes_{R}1_{R_{\mathfrak{m}}})(m\otimes_{R}1_{R_{\mathfrak{m}}})=(a\otimes_{R}1_{R_{\mathfrak{m}}})(1\otimes_{R}m)$) and thus $b=c\otimes_{R}\dfrac{1}{s}=(cs\otimes_{R}1_{R_{\mathfrak{m}}})\dfrac{1}{s^{2}}\in(A\otimes_{R}R_{\mathfrak{m}})\mathfrak{m}_{2}$.
The statement about the maximality of $\alpha_{2}$ follows from the fact that the homeomorphism in [15, Proposition 3.3.10] clearly preserves inclusions. ∎
Lemma 2.16.
Let $\mathfrak{m}$ be a maximal ideal of $R$ and let $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$. Then there is $a\in\operatorname{Sym}(A,\sigma)$ such that $a\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$ and $a\not\in A\mathfrak{m}$.
Proof.
We assume that the conclusion does not hold, so that property (2.2) of Lemma 2.15 holds. We proceed in four steps:
Step 1: Take $a\in\operatorname{Sym}(A,\sigma)$ such that $a\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$, cf. Lemma 2.13. Then $a\in A\mathfrak{m}$ by property (2.2). Furthermore, for every $r\in\operatorname{Supp}\alpha$ we have $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a+r\rangle_{\sigma}=\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}$, cf. Definition 2.1 (since the first step in the computation is scalar extension to $\mathrm{qf}(R/\operatorname{Supp}\alpha)$). Thus, by hypothesis, $a+r\in A\mathfrak{m}$. Since $a\in A\mathfrak{m}$, we get that $\operatorname{Supp}\alpha\subseteq A\mathfrak{m}\cap R=\mathfrak{m}$ (cf. [9, Corollary 7.1.2(1)] for the equality).
Step 2: We first apply Lemma 2.15(1) and get that we can assume that $\operatorname{Supp}\alpha=\{0\}$, and in particular that $R$ is a domain. It is then possible to apply Lemma 2.15(2) and we can also assume that $R$ is a local domain with maximal ideal $\mathfrak{m}$.
Step 3: Since $R$ is a local domain and $\operatorname{Supp}\alpha=\{0\}$, the following holds:
For every $\frac{r_{1}}{s_{1}}\in\mathrm{qf}(R)$ with $r_{1},s_{1}\in\alpha\setminus\{0\}$, there exists $r\in\alpha\cap R^{\times}$ such that $r\leq_{\alpha}\frac{r_{1}}{s_{1}}$.
Proof of this claim: By the description of $\alpha$ in Theorem 1.19, there is $r_{1}^{\prime}\in\alpha\cap R^{\times}$ such that $r_{1}^{\prime}\leq_{\alpha}r_{1}$, so that $\frac{r_{1}^{\prime}}{s_{1}}\leq_{\alpha}\frac{r_{1}}{s_{1}}$. Observe that if there is $s^{\prime}_{1}\in\alpha\cap R^{\times}$ such that $s^{\prime}_{1}\geq_{\alpha}s_{1}$, then $\frac{r_{1}^{\prime}}{s^{\prime}_{1}}\leq_{\alpha}\frac{r_{1}}{s_{1}}$, and we can take $r:=r^{\prime}_{1}{s^{\prime}_{1}}^{-1}$.
[…]
Step 4: We work in the central simple algebra with involution $(A(\alpha),\sigma(\alpha))=(A\otimes_{R}\mathrm{qf}(R),\sigma\otimes\operatorname{id})$, and denote by $\bar{\alpha}$ the ordering induced by $\alpha$ on $\mathrm{qf}(R)$. By Lemma 2.14(2), there is $\frac{r_{1}}{s_{1}}\in\bar{\alpha}\setminus\{0\}$ such that $\langle a\otimes 1-\dfrac{r_{2}}{s_{2}}\rangle_{\sigma(\alpha)}$ has maximal signature at $\bar{\alpha}$ for every $\frac{r_{2}}{s_{2}}\in\bar{\alpha}$ such that $\frac{r_{2}}{s_{2}}\leq_{\bar{\alpha}}\frac{r_{1}}{s_{1}}$. In other words, $a\otimes 1-\frac{r_{1}}{r_{2}}\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$.
By Step 3, there is $r\in\alpha\cap R^{\times}$ such that $r\leq_{\bar{\alpha}}\frac{r_{1}}{s_{1}}$. In particular we have $(a-r)\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. Therefore, $a-r\in A\mathfrak{m}$ by property (2.2) and thus $r\in A\mathfrak{m}$. But this is impossible since $A\mathfrak{m}$ is a proper ideal and $r$ is invertible. ∎
Lemma 2.17.
Assume that $R$ is semilocal, let $\mathfrak{m}$ be a maximal ideal of $R$, and let $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$. Then there is $b_{\mathfrak{m}}\in\mathscr{S}_{\alpha}(A,\sigma)$ such that $b_{\mathfrak{m}}+A\mathfrak{m}\in(A/A\mathfrak{m})^{\times}$.
Proof.
Let $a\in\operatorname{Sym}(A,\sigma)$ be such that $a\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$ and $\pi_{A\mathfrak{m}}(a)\not=0$, cf. Lemma 2.16. Then $D_{(A/A\mathfrak{m},\bar{\sigma})}(k\times\langle\pi_{A\mathfrak{m}}(a)\rangle_{\bar{\sigma}})\subseteq\mathscr{S}^{\prime}_{\mathfrak{m}}$ for all $k\in\mathbb{N}$ by definition of $\mathscr{S}^{\prime}_{\mathfrak{m}}$. Since $(A/A\mathfrak{m},\bar{\sigma})$ is a central simple algebra with involution, [4, Lemma 2.4] applies and there is $\ell\in\mathbb{N}$ such that ×$\ell\times\langle\pi_{A\mathfrak{m}}(a)\rangle_{\bar{\sigma}}$ represents an invertible element $b^{\prime}$. Since $b^{\prime}\in\mathscr{S}^{\prime}_{\mathfrak{m}}$ and $\pi_{A\mathfrak{m}}$ is surjective, we have
$b^{\prime}=\sum_{j=1}^{\ell}\bar{\sigma}(\pi_{A\mathfrak{m}}(x_{j}))\pi_{A\mathfrak{m}}(a)\pi_{A\mathfrak{m}}(x_{j}),$
with $x_{j}\in A$. Therefore, we take $b_{\mathfrak{m}}=\sum_{j=1}^{\ell}\sigma(x_{j})ax_{j}$. ∎
Proposition 2.18.
Assume that $R$ is semilocal and let $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$. Then there are invertible elements in $\mathscr{S}_{\alpha}(A,\sigma)$. Furthermore, every invertible element $a\in\mathscr{S}_{\alpha}(A,\sigma)$ satisfies $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}=m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$.
Proof.
Let $\mathfrak{m}_{1},\ldots,\mathfrak{m}_{\ell}$ be the maximal ideals of $R$. Observe that for each $i\in\{1,\ldots,\ell\}$ there is $b_{i}\in\mathscr{S}_{\alpha}(A,\sigma)$ such that $b_{i}+A\mathfrak{m}_{i}\in(A/A\mathfrak{m}_{i})^{\times}$ by Lemma 2.17. By the Chinese remainder theorem, the canonical map $\xi\colon R\rightarrow R/\mathfrak{m}_{1}\times\cdots\times R/\mathfrak{m}_{\ell}$ is surjective. In particular there are $r_{1},\ldots,r_{\ell}\in R$ such that $\xi(r_{i})=(0,\ldots,0,1,0,\ldots,0)$, where the coordinate $1$ is the one corresponding to the quotient $R/{\mathfrak{m}_{i}}$. Define
$b:=\sigma(r_{1})b_{1}r_{1}+\cdots+\sigma(r_{\ell})b_{\ell}r_{\ell}.$
Observe that $b\in\mathscr{S}_{\alpha}(A,\sigma)$. We check that $b$ is invertible. By Proposition 1.10, it suffices to show that $b+A\mathfrak{m}$ is invertible in $A/A\mathfrak{m}$ for every maximal ideal $\mathfrak{m}$ of $R$. Consider such an ideal $\mathfrak{m}_{i}$. By definition of $r_{i}$ we have $r_{i}+\mathfrak{m}_{i}=1+\mathfrak{m}_{i}$ and $r_{j}+\mathfrak{m}_{i}=0+\mathfrak{m}_{i}$ for all $j\not=i$. Therefore $b+A\mathfrak{m}_{i}=b_{i}+A\mathfrak{m}_{i}$, which is invertible in $A/A\mathfrak{m}_{i}$.
We show that if $a\in\mathscr{S}_{\alpha}(A,\sigma)$ is invertible, then $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}=m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. Since $a\in\mathscr{S}_{\alpha}(A,\sigma)$, there are $a_{1},\ldots,a_{k}\in\operatorname{Sym}(A,\sigma)$ such that $a\in D_{(A,\sigma)}\langle a_{1},\ldots,a_{k}\rangle_{\sigma}$ and $a_{1}\otimes 1,\ldots,a_{k}\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. Since $a$ is invertible, a standard argument gives $\langle a_{1},\ldots,a_{k}\rangle_{\sigma}\simeq\langle a\rangle_{\sigma}\perp h$ for some hermitian form $h$ over $(A,\sigma)$. Extending the scalars to $\kappa(\alpha)$, we obtain $\langle a\otimes 1\rangle_{\sigma(\alpha)}\perp h\otimes\kappa(\alpha)\simeq\langle a_{1}\otimes 1,\ldots,a_{k}\otimes 1\rangle_{\sigma(\alpha)}$ over the central simple algebra with involution $(A(\alpha),\sigma(\alpha))$. Observe that $a_{1}\otimes 1,\ldots,a_{k}\otimes 1$ are invertible in $A(\alpha)$, and thus that $h\otimes\kappa(\alpha)$ is nonsingular. By [4, Lemma 2.2] there is $\ell\in\mathbb{N}$ such that ×$\ell\times(h\otimes\kappa(\alpha))\simeq\langle c_{1},\ldots,c_{t}\rangle_{\sigma(\alpha)}$ for some $c_{1},\ldots,c_{t}\in\operatorname{Sym}(A(\alpha)^{\times},\sigma(\alpha))$ (they are invertible since $h\otimes\kappa(\alpha)$ is nonsingular). Therefore,
×$\ell\times\langle a\otimes 1\rangle_{\sigma(\alpha)}\perp\langle c_{1},\ldots,c_{t}\rangle_{\sigma(\alpha)}\simeq\ell\times\langle a_{1}\otimes 1,\ldots,a_{k}\otimes 1\rangle_{\sigma(\alpha)}.$
Note that both forms are diagonal over $(A(\alpha),\sigma(\alpha))$, so that $\ell+t=\ell k$ for dimension reasons. Since $a_{1}\otimes 1,\ldots,a_{k}\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$, the form on the right-hand side has the maximal signature that can be obtained by a nonsingular diagonal form of dimension $\ell k$ over $(A(\alpha),\sigma(\alpha))$, namely $\ell k\cdot m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. It is therefore the same for the form on the left-hand side, which implies that $a\otimes 1$ (and every $c_{i}$) belongs to $M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$, i.e., $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}=m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$. ∎
Since $m_{\alpha}(A,\sigma)\leq m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$ by the definition of signatures, the following corollary is an immediate consequence of Proposition 2.18:
Corollary 2.19.
Assume that $R$ is semilocal and that $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$. Then
  1. (1)
    $m_{\alpha}(A,\sigma)=m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$.
  2. (2)
    If $a\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$, then $a\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$.
Note that if $\alpha\not\in\operatorname{Nil}[A,\sigma]$, then $m_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))=n_{\bar{\alpha}}$.
Proof.
The final statement is the only one that still requires a proof and follows from [4, Proposition 6.7] since $\bar{\alpha}\not\in\operatorname{Nil}[A(\alpha),\sigma(\alpha)]$ by Lemma 2.7. ∎
As already mentioned in Remark 2.2, we have $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}\leq n_{\bar{\alpha}}$ for $a\in\operatorname{Sym}(A,\sigma)$ by [3, Proposition 4.4]. It immediately follows from Corollary 2.19 that the definition of $m_{\alpha}(A,\sigma)$ could include non-invertible elements when $R$ is semilocal:
Corollary 2.20.
Assume that $R$ is semilocal and that $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$. Then
$m_{\alpha}(A,\sigma)=\max\{\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle a\rangle_{\sigma}\mid a\in\operatorname{Sym}(A,\sigma)\}.$
Remark 2.21.
While an element of maximal signature in a central simple algebra with involution over a field is necessarily invertible (cf. Lemma 2.14), this may not be so for Azumaya algebras with involution, even already in the ring case. For example, let $(A,\sigma)=(\mathbb{Z}_{3\mathbb{Z}},\operatorname{id})$ (in particular, hermitian forms over $(A,\sigma)$ are just bilinear forms over $\mathbb{Z}_{3\mathbb{Z}}$ and their signatures are the usual Sylvester signatures). The ring $\mathbb{Z}_{3\mathbb{Z}}$ has a unique ordering $\alpha_{0}$, and $\operatorname{sign}_{\alpha_{0}}\langle 3\rangle=1=m_{\alpha_{0}}(A,\sigma)$.
3. The involution trace pairing
Let $R$ be a commutative ring (with $2\in R^{\times}$), let $(A,\sigma)$ be an Azumaya algebra with involution over $R$, let $S=Z(A)$ and let $\iota:=\sigma|_{S}$. Note that $A$ is Azumaya over $S$, but not necessarily Azumaya over $R$.
3.1. The involution trace form
We consider the reduced trace of $A$, $\operatorname{Trd}_{A}\colon A\to S$, cf. [18, IV, §2], and recall that it is additive and $S$-linear. Furthermore, $\operatorname{Trd}_{A}$ commutes with scalar extensions of $S$ since its computation does not depend on the choice of splitting ring, cf. [18, IV, Proposition 2.1]. In fact, $\operatorname{Trd}_{A}$ also commutes with scalar extensions of $R$ (the case of interest to us) as the following computation shows:
Lemma 3.1.
Let $R^{\prime}$ be a commutative ring that contains $R$. Then for all $a\in A$,
$\operatorname{Trd}_{A}(a)\otimes_{R}1_{R^{\prime}}=\operatorname{Trd}_{A\otimes_{R}R^{\prime}}(a\otimes_{R}1_{R^{\prime}}).$
Proof.
Observe that $A\otimes_{R}R^{\prime}\cong A\otimes_{S}(S\otimes_{R}R^{\prime})$, via $a\otimes_{R}r^{\prime}\mapsto a\otimes_{S}1\otimes_{R}r^{\prime}$. It follows that
$\operatorname{Trd}_{A}(a)\otimes_{R}1_{R^{\prime}}$ $=\operatorname{Trd}_{A}(a)\otimes_{S}1_{S}\otimes_{R}1_{R^{\prime}}$
$=\operatorname{Trd}_{A\otimes_{S}S\otimes_{R}R^{\prime}}(a\otimes_{S}1_{S}\otimes_{R}1_{R^{\prime}})$
$=\operatorname{Trd}_{A\otimes_{R}R^{\prime}}(a\otimes_{R}1_{R^{\prime}}).\qed$
Lemma 3.2.
For all $a\in A$ we have $\operatorname{Trd}_{A}(\sigma(a))=\iota(\operatorname{Trd}_{A}(a))$.
Proof.
If $\iota=\mathrm{id}_{S}$, i.e., $S=R$, the statement follows from [16, III, (8.1.1)]. We assume that $\iota\not=\mathrm{id}_{S}$, and first observe that
$\operatorname{Trd}_{A}$ $(\sigma(a))=\iota(\operatorname{Trd}_{A}(a))\Leftrightarrow\forall\mathfrak{p}\in\operatorname{Spec}R\ \operatorname{Trd}_{A}(\sigma(a))\otimes_{R}1_{R_{\mathfrak{p}}}=\iota(\operatorname{Trd}_{A}(a))\otimes_{R}1_{R_{\mathfrak{p}}}$
$\Leftrightarrow\forall\mathfrak{p}\in\operatorname{Spec}R\ \operatorname{Trd}_{A\otimes_{R}R_{\mathfrak{p}}}(\sigma(a)\otimes 1_{R_{\mathfrak{p}}})=(\iota\otimes\operatorname{id})(\operatorname{Trd}_{A}(a)\otimes 1_{R_{\mathfrak{p}}})$
$\Leftrightarrow\forall\mathfrak{p}\in\operatorname{Spec}R\ \operatorname{Trd}_{A\otimes_{R}R_{\mathfrak{p}}}((\sigma\otimes\operatorname{id})(a\otimes 1_{R_{\mathfrak{p}}}))=(\iota\otimes\operatorname{id})(\operatorname{Trd}_{A\otimes_{R}R_{\mathfrak{p}}}(a\otimes 1_{R_{\mathfrak{p}}})).$
Therefore, it suffices to prove the result for $(A\otimes_{R}R_{\mathfrak{p}},\sigma\otimes\operatorname{id})$, and in particular it suffices to prove the statement of the lemma under the extra hypothesis that $R$ is local. In this case, the arguments in [17, (2.15) and (2.16)] hold mutatis mutandis for the Azumaya algebra with involution $(A,\sigma)$ over $R$. More precisely, since $R$ is local, there is $\lambda\in S$ such that $\lambda^{2}\in R^{\times}$, $\iota(\lambda)=-\lambda$ and $S=R\oplus\lambda R$, cf. Section 1.2. Then [17, (2.15)] holds for $(A,\sigma)$, i.e., the map
$(A\otimes_{R}S,\sigma\otimes\operatorname{id})\to(A\times A^{\mathrm{op}},\varepsilon),\ a\otimes s\mapsto(as,(\sigma(a)s)^{\mathrm{op}}),$
where $\varepsilon$ denotes the exchange involution, is an isomorphism of $S$-algebras with involution (replacing the element $\alpha$ in the proof of [17, (2.15)] by $\lambda$, and observing that $\lambda-\iota(\lambda)=2\lambda\in S^{\times}$). The claimed equality becomes straightforward to verify after application of this isomorphism since the reduced trace is invariant under scalar extension. ∎
The involution trace form of $(A,\sigma)$ is the form
$T_{\sigma}\colon A\times A\to S,\ (x,y)\mapsto\operatorname{Trd}_{A}(\sigma(x)y).$
By Lemma 3.2, $T_{\sigma}$ is symmetric bilinear over $R$ if $S=R$ and hermitian over $(S,\iota)$ otherwise.
Lemma 3.3.
The form $T_{\sigma}$ is nonsingular, i.e., $(A,T_{\sigma})\in\mathfrak{H}(S,\iota)$.
Proof.
Since $S$ is a finitely generated $R$-module (cf. Definition 1.4 and Proposition 1.3), the form $T_{\sigma}$ is nonsingular if and only if the form $T_{\sigma}\otimes_{R}R/\mathfrak{m}$ is nonsingular for all maximal ideals $\mathfrak{m}$ of $R$, cf. [16, I, Lemma 7.1.3]. Since $\operatorname{Trd}_{A}$ commutes with scalar extension, $T_{\sigma}\otimes_{R}R/\mathfrak{m}$ is isometric to the involution trace form of the central simple algebra with involution $(A\otimes_{R}R/\mathfrak{m},\sigma\otimes\operatorname{id})$, which is nonsingular, cf. [17, §11]. ∎
Lemma 3.4.
The sandwich map $\mathrm{sw}$ (cf. (1.3)) induces an isomorphism
$(A\otimes_{S}A^{\mathrm{op}},\sigma\otimes\sigma^{\mathrm{op}})\cong(\operatorname{End}_{S}(A),\operatorname{ad}_{T_{\sigma}})$
of Azumaya algebras with involution over $R$.
Proof.
By (1.3), the sandwich map is an isomorphism of $S$-algebras. We first show that it respects the involutions, i.e., $\operatorname{ad}_{T_{\sigma}}(\mathrm{sw}(a\otimes b^{\mathrm{op}}))=\mathrm{sw}(\sigma(a)\otimes\sigma^{\mathrm{op}}(b^{\mathrm{op}}))$ for all $a,b\in A$. With reference to (1.1) this follows from the straightforward computation
$T_{\sigma}(x,\mathrm{sw}(\sigma(a)\otimes\sigma^{\mathrm{op}}(b^{\mathrm{op}}))(y))$ $=\operatorname{Trd}_{A}(\sigma(x)\sigma(a)y\sigma(b))=\operatorname{Trd}_{A}(\sigma(b)\sigma(x)\sigma(a)y)$
$=\operatorname{Trd}_{A}(\sigma(axb)y)=T_{\sigma}(axb,y)$
$=T_{\sigma}(\mathrm{sw}(a\otimes b^{\mathrm{op}})(x),y)$
which holds for all $x,y,a,b\in A$.
Finally, $(\operatorname{End}_{S}(A),\operatorname{ad}_{T_{\sigma}})$ is an Azumaya algebra with involution over $R$. Indeed, by Proposition 1.3, $S$ is finite étale over $R$ and $A$ is Azumaya over $S$. Hence, $\operatorname{End}_{S}(A)$ is Azumaya over $S$ and in particular has centre $S$. Therefore, by Proposition 1.3 $\operatorname{End}_{S}(A)$ is projective and separable over $R$. Clearly, $\operatorname{ad}_{T_{\sigma}}$ is $R$-linear. Thus we just have to show that if $s\in S\setminus R$, then $\operatorname{ad}_{T_{\sigma}}(s\cdot\operatorname{id}_{A})\not=s\cdot\operatorname{id}_{A}$. This can be checked by showing that if $\operatorname{ad}_{T_{\sigma}}(s\cdot\operatorname{id}_{A})=s\cdot\operatorname{id}_{A}$ for some $s\in S$, then $\sigma(s)=s$, using the nonsingularity of $T_{\sigma}$ and the definition of $\operatorname{ad}_{T_{\sigma}}$ in a similar fashion to the computation above. ∎
3.2. The Goldman element
We can view $\operatorname{Trd}_{A}$ as an element of $\operatorname{End}_{S}(A)$. By the definition of the sandwich isomorphism (1.3), there is a unique element $g_{A}=\sum_{i}x_{i}\otimes y^{\mathrm{op}}_{i}$ in $A\otimes_{S}A^{\mathrm{op}}$ such that
(3.1) $\mathrm{sw}(g_{A})(a)=\sum_{i}x_{i}ay_{i}=\operatorname{Trd}_{A}(a)\quad\text{for all }a\in A.$
The element $g_{A}$ is called the Goldman element of $A$.
Lemma 3.5.
The Goldman element $g_{A}$ satisfies
(3.2) $(\sigma\otimes\sigma^{\mathrm{op}})(g_{A})=g_{A}.$
Proof.
By Lemma 3.4 it suffices to show that $\operatorname{ad}_{T_{\sigma}}(\operatorname{Trd}_{A})=\operatorname{Trd}_{A}$ in $\operatorname{End}_{S}(A)$ in order to prove the claim. Consider (1.1) with $h=T_{\sigma}$ and $f=\operatorname{Trd}_{A}$. Using the properties of $T_{\sigma}$ and $\operatorname{Trd}_{A}$ we have
$T_{\sigma}(x,\operatorname{ad}_{T_{\sigma}}(\operatorname{Trd}_{A})(y))$ $=T_{\sigma}(\operatorname{Trd}_{A}(x),y)=\iota(\operatorname{Trd}_{A}(x))T_{\sigma}(1,y)$
$=\operatorname{Trd}_{A}(\sigma(x))T_{\sigma}(1,y)=T_{\sigma}(x,1)\operatorname{Trd}_{A}(y)$
$=T_{\sigma}(x,\operatorname{Trd}_{A}(y))$
for all $x,y\in A$. Since $T_{\sigma}$ is nonsingular, the claim follows. ∎
We usually think of $g_{A}$ as an element of $A\otimes_{S}A$ via the canonical $S$-module isomorphism $A\otimes_{S}A\to A\otimes_{S}A^{\mathrm{op}},a\otimes b\mapsto a\otimes b^{\mathrm{op}}$ and write $g_{A}=\sum_{i}x_{i}\otimes y_{i}$, cf. [18, p. 112]. Since $\sigma\otimes\sigma$ and $\sigma\otimes\sigma^{\mathrm{op}}$ correspond to each other as additive maps under this isomorphism, Lemma 3.5 yields
(3.3) $(\sigma\otimes\sigma)(g_{A})=g_{A}.$
in $A\otimes_{S}A$. Furthermore, we have
(3.4) $g_{A}^{2}=1$
and
(3.5) $g_{A}(a\otimes b)=(b\otimes a)g_{A}\quad\text{for all }a,b\in A,$
cf. [18, IV, Proposition 4.1].
3.3. Module actions
We define $\prescript{\iota}{}{A}$ to be the $S$-algebra given by the ring $A$ equipped with the following left action by $S$:
$S\times A\rightarrow A,\quad(s,a)\mapsto s\prescript{\iota}{}{\!}\!\cdot a:=a\iota(s).$
We denote this action of $S$ on $A$ with the symbol $\prescript{\iota}{}{\!}\!\cdot{}$ in order to distinguish it from the product in $A$ of elements of $S$ and $A$. Note that $(\prescript{\iota}{}{A},\sigma)$ is an $(S,\iota)$-algebra with involution as presented in Section 1.1, and that if $(S,\iota)=(R,\operatorname{id})$, then $\prescript{\iota}{}{A}=A$.
We can view $A$ as a left $A\otimes_{S}\prescript{\iota}{}{A}$-module via the twisted sandwich action:
(3.6) $a\otimes b\cdot_{\mathrm{ts}}x:=ax\sigma(b)$
for all $a,b,x\in A$ (it is necessary to use $\prescript{\iota}{}{A}$ in the tensor product instead of $A$, in order for the action to be well-defined, which is the motivation for introducing $\prescript{\iota}{}{A}$).
Lemma 3.6.
The twisted sandwich map
$A\otimes_{S}\prescript{\iota}{}{A}\to\operatorname{End}_{S}(A),a\otimes b\mapsto[x\mapsto ax\sigma(b)]$
induces an isomorphism
$(A\otimes_{S}\prescript{\iota}{}{A},\sigma\otimes\sigma)\cong(\operatorname{End}_{S}(A),\operatorname{ad}_{T_{\sigma}})$
of Azumaya algebras with involution over $R$.
Proof.
The map $\sigma\colon A^{\mathrm{op}}\to\prescript{\iota}{}{A}$ is an isomorphism of $S$-algebras and yields an isomorphism $(A\otimes_{S}A^{\mathrm{op}},\sigma\otimes\sigma^{\mathrm{op}})\cong(A\otimes_{S}\prescript{\iota}{}{A},\sigma\otimes\sigma)$ of Azumaya algebras with involution over $R$. The result then follows from Lemma 3.4. ∎
For a right $A$-module $M$ we denote by $\prescript{\iota}{}{M}$ the right $\prescript{\iota}{}{A}$-module with the same elements as $M$, with multiplication $\prescript{\iota}{}{M}\times\prescript{\iota}{}{A}\to\prescript{\iota}{}{M}$ the same as the multiplication $M\times A\to M$ and with left action by $S$ given by
$S\times\prescript{\iota}{}{M}\rightarrow\prescript{\iota}{}{M},\quad(s,m)\mapsto s\prescript{\iota}{}{\!}\!\cdot m:=m\iota(s).$
If $(M,h)\in\mathfrak{Herm}^{\varepsilon}(A,\sigma)$, then $h$ is $\iota(\varepsilon)$-hermitian on $\prescript{\iota}{}{M}$, and we denote it by $\prescript{\iota}{}{h}$.
If $M_{1}$ and $M_{2}$ are right $A$-modules a direct verification shows that $M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2}$ is a right $A\otimes_{S}\prescript{\iota}{}{A}$-module with multiplication induced by
$(M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\times(A\otimes_{S}\prescript{\iota}{}{A})\rightarrow M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2},\quad(m_{1}\otimes m_{2})\cdot(a\otimes b):=m_{1}a\otimes m_{2}b.$
3.4. The involution trace pairing
Let $(M_{1},h_{1})\in\mathfrak{Herm}^{\varepsilon_{1}}(A,\sigma)$ and $(M_{2},h_{2})\in\mathfrak{Herm}^{\varepsilon_{2}}(A,\sigma)$, and consider the involution trace form $(A,T_{\sigma})\in\mathfrak{H}(S,\iota)$. Using the twisted sandwich action of $A\otimes_{S}\prescript{\iota}{}{A}$ on $A$ (cf. (3.6)), we can define the $S$-module $(M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A$, which carries the form $T_{\sigma}\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2})\in\mathfrak{Herm}^{\varepsilon_{1}\iota(\varepsilon_{2})}(S,\iota)$, where $\bullet$ denotes the product of forms from [16, I, (8.1), (8.2)]. In other words,
$T_{\sigma}\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2})(m_{1}\otimes m_{2}\otimes a,m_{1}^{\prime}\otimes m_{2}^{\prime}\otimes a^{\prime}):=T_{\sigma}\bigl{(}a,h_{1}(m_{1},m_{1}^{\prime})\otimes\prescript{\iota}{}{h}_{2}(m_{2},m_{2}^{\prime})\cdot_{\mathrm{ts}}a^{\prime}\bigr{)}$
for all $m_{1},m_{1}^{\prime}\in M_{1}$, $m_{2},m_{2}^{\prime}\in\prescript{\iota}{}{M}_{2}$ and $a,a^{\prime}\in A$. (Note that we do not indicate all parentheses in long tensor products of elements, in order not to overload the notation.)
For this product to be well-defined, $A$ needs to be an $(A\otimes_{S}\prescript{\iota}{}{A})$-$S$ bimodule and the form $(A,T_{\sigma})$ needs to “admit” $A\otimes_{S}\prescript{\iota}{}{A}$, which means that the equality
$T_{\sigma}(\sigma(x)\otimes\sigma(y)\cdot_{\mathrm{ts}}a,b)=T_{\sigma}(a,x\otimes y\cdot_{\mathrm{ts}}b)$
must hold for all $x,y,a,b\in A$, but this follows from Lemma 3.6.
In this way we obtain the pairing
(3.7) $\begin{split}*\colon&\mathfrak{Herm}^{\varepsilon_{1}}(A,\sigma)\times\mathfrak{Herm}^{\varepsilon_{2}}(A,\sigma)\to\mathfrak{Herm}^{\varepsilon_{1}\iota(\varepsilon_{2})}(S,\iota),\\ (M_{1},h_{1})&*(M_{2},h_{2}):=\bigl{(}(M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A,T_{\sigma}\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2})\bigr{)}.\end{split}$
Expanding the definition of the pairing and simply writing $h_{1}*h_{2}$, we see that
$h_{1}*h_{2}(m_{1}\otimes m_{2}\otimes a,m_{1}^{\prime}\otimes m_{2}^{\prime}\otimes a^{\prime})$ $=T_{\sigma}\bigl{(}a,h_{1}(m_{1},m_{1}^{\prime})\otimes\prescript{\iota}{}{h}_{2}(m_{2},m_{2}^{\prime})\cdot_{\mathrm{ts}}a^{\prime}\bigr{)}$
(3.8) $=\operatorname{Trd}_{A}\bigl{(}\sigma(a)h_{1}(m_{1},m_{1}^{\prime})a^{\prime}\sigma(h_{2}(m_{2},m_{2}^{\prime}))\bigr{)}$
$=\operatorname{Trd}_{A}\bigl{(}h_{1}(m_{1}a,m_{1}^{\prime}a^{\prime})\sigma(h_{2}(m_{2},m_{2}^{\prime}))\bigr{)}.$
Lemma 3.7.
The Azumaya algebras with involution $(A\otimes_{S}\prescript{\iota}{}{A},\sigma\otimes\sigma)$ and $(S,\iota)$ are Morita equivalent via
$\mathfrak{Herm}^{\varepsilon}(A\otimes_{S}\prescript{\iota}{}{A},\sigma\otimes\sigma)\to\mathfrak{Herm}^{\varepsilon}(S,\iota),\ \varphi\mapsto T_{\sigma}\bullet\varphi.$
Proof.
This follows from Lemma 3.6 and [16, I, Theorem 9.3.5]. ∎
Corollary 3.8.
  1. (1)
    The pairing * preserves orthogonal sums in each component.
  2. (2)
    If $h_{1}$ and $h_{2}$ are nonsingular, then $h_{1}*h_{2}$ is nonsingular.
  3. (3)
    If $h_{1}\simeq h_{1}^{\prime}$ and $h_{2}\simeq h_{2}^{\prime}$, then $h_{1}*h_{2}\simeq h_{1}^{\prime}*h_{2}^{\prime}$.
Proof.
Let $h_{i}\in\mathfrak{Herm}^{\varepsilon_{i}}(A,\sigma)$ for $i=1,2$, then $h_{1}*h_{2}=T_{\sigma}\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2})$ and the three statements follow from Lemma 3.7, [16, I, Theorem 9.3.5] and standard properties of the tensor product of forms. ∎
Theorem 3.9.
Let $(M_{i},h_{i})\in\mathfrak{Herm}^{\varepsilon_{i}}(A,\sigma)$ for $i=1,2,3$. Then
$(h_{1}*h_{2})\otimes_{S}h_{3}\simeq(h_{3}*h_{2})\otimes_{S}h_{1}.$
Proof.
We are grateful to the first referee for suggesting this proof, which is significantly shorter and more conceptual than our original one.
Consider the $\varepsilon_{1}\iota(\varepsilon_{2})\varepsilon_{3}$-hermitian forms $(M,h):=(M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2}\otimes_{S}M_{3},h_{1}\otimes\prescript{\iota}{}{h}_{2}\otimes h_{3})$ and $(M^{\prime},h^{\prime}):=(M_{3}\otimes_{S}\prescript{\iota}{}{M}_{2}\otimes_{S}M_{1},h_{3}\otimes\prescript{\iota}{}{h}_{2}\otimes h_{1})$ over $(A\otimes_{S}\prescript{\iota}{}{A}\otimes_{S}A,\sigma^{\otimes 3})$. Let $g^{\prime}_{A}\in A\otimes_{S}\prescript{\iota}{}{A}\otimes_{S}A$ be the image of the Goldman element $g_{A}\in A\otimes_{S}A$ under the natural map $a\otimes b\mapsto a\otimes 1\otimes b$. Using the properties of $g^{\prime}_{A}$, induced by those of $g_{A}$ (cf. Section 3.2), the following computation shows that $(M,h)$ and $(M^{\prime},h^{\prime})$ are isometric via the isomorphism of right $A\otimes_{S}\prescript{\iota}{}{A}\otimes_{S}A$-modules $M\to M^{\prime}$, defined by $m_{1}\otimes m_{2}\otimes m_{3}\mapsto(m_{3}\otimes m_{2}\otimes m_{1})g^{\prime}_{A}$:
$h^{\prime}((m_{3}\otimes m_{2}\otimes m_{1})g^{\prime}_{A},$ $(m^{\prime}_{3}\otimes m^{\prime}_{2}\otimes m^{\prime}_{1})g^{\prime}_{A})$
$=\sigma^{\otimes 3}(g^{\prime}_{A})h^{\prime}(m_{3}\otimes m_{2}\otimes m_{1},m^{\prime}_{3}\otimes m^{\prime}_{2}\otimes m^{\prime}_{1})g^{\prime}_{A}$
$=g^{\prime}_{A}(h_{3}(m_{3},m^{\prime}_{3})\otimes\prescript{\iota}{}{h}_{2}(m_{2},m^{\prime}_{2})\otimes h_{1}(m_{1},m^{\prime}_{1}))g^{\prime}_{A}$
$=h_{1}(m_{1},m^{\prime}_{1})\otimes\prescript{\iota}{}{h}_{2}(m_{2},m^{\prime}_{2})\otimes h_{3}(m_{3},m^{\prime}_{3})$
$=h(m_{1}\otimes m_{2}\otimes m_{3},m^{\prime}_{1}\otimes m^{\prime}_{2}\otimes m^{\prime}_{3}).$
Since $(A,\sigma)\cong(\operatorname{End}_{A}(A),\operatorname{ad}_{\langle 1\rangle_{\sigma}})$, it follows from Lemmas 3.6 and 1.2 that
$(A\otimes_{S}\prescript{\iota}{}{A}\otimes_{S}A,\sigma^{\otimes 3})$ $\cong(\operatorname{End}_{S}(A)\otimes_{S}A,\operatorname{ad}_{T_{\sigma}}\otimes\sigma)$
$\cong(\operatorname{End}_{S}(A)\otimes_{S}\operatorname{End}_{A}(A),\operatorname{ad}_{T_{\sigma}}\otimes\operatorname{ad}_{\langle 1\rangle_{\sigma}})$
$\cong(\operatorname{End}_{S\otimes_{S}A}(A\otimes_{S}A),\operatorname{ad}_{T_{\sigma}\otimes\langle 1\rangle_{\sigma}}),$
which yields the Morita equivalence
(3.9) $\mathfrak{Herm}^{\varepsilon}(A\otimes_{S}\prescript{\iota}{}{A}\otimes_{S}A,\sigma^{\otimes 3})\to\mathfrak{Herm}^{\varepsilon}(S\otimes_{S}A,\iota\otimes\sigma),\ \varphi\mapsto(T_{\sigma}\otimes\langle 1\rangle_{\sigma})\bullet\varphi$
by [16, I, Theorem 9.3.5]. The isomorphism of right $A\otimes_{S}A$-modules
$((M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{S}M_{3})\otimes_{(A\otimes_{S}\prescript{\iota}{}{A})\otimes_{S}A}(A\otimes_{S}A)\rightarrow((M_{1}\otimes_{s}\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A)\otimes_{S}(M_{3}\otimes_{A}A),$
$m_{1}\otimes m_{2}\otimes m_{3}\otimes a\otimes b\mapsto m_{1}\otimes m_{2}\otimes a\otimes m_{3}\otimes b,$
followed by the isomorphism of right $A$-modules $M_{3}\otimes_{A}A\rightarrow M_{3}$, $m_{3}\otimes b\mapsto m_{3}b$, then yield the isometries
$(T_{\sigma}\otimes\langle 1\rangle_{\sigma})\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2}\otimes h_{3})\simeq T_{\sigma}\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2})\otimes\langle 1\rangle_{\sigma}\bullet h_{3}\simeq(h_{1}*h_{2})\otimes h_{3}.$
A similar argument shows that $(T_{\sigma}\otimes\langle 1\rangle_{\sigma})\bullet(h_{3}\otimes\prescript{\iota}{}{h}_{2}\otimes h_{1})\simeq(h_{3}*h_{2})\otimes h_{1}$. The result then follows from the isometry $(M,h)\simeq(M^{\prime},h^{\prime})$ since (3.9) preserves isometries. ∎
Remark 3.10.
The pairing * was introduced and studied in detail for $\varepsilon$-hermitian forms over central simple algebras with involution by Garrel in [10]. (A similar construction for quaternion algebras had already been considered by Lewis [20], using the norm form instead of the involution trace form of the quaternion conjugation, cf. [10, Remark 4.4].) In our presentation we stayed close to Garrel’s approach via hermitian Morita theory. We are grateful to the second referee for suggesting an alternative approach via the $S$-linear isomorphism
$M_{1}\otimes_{A}\prescript{\sigma}{}{M}_{2}\to(M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A,\ m_{1}\otimes m_{2}\mapsto(m_{1}\otimes m_{2})\otimes 1$
(where $\prescript{\sigma}{}{M}_{2}$ is the left $A$-module obtained by twisting the right $A$-module structure of $M_{2}$ by $\sigma$) with inverse
$(M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A\to M_{1}\otimes_{A}\prescript{\sigma}{}{M}_{2},(m_{1}\otimes m_{2})\otimes a\mapsto m_{1}a\otimes m_{2}=m_{1}\otimes m_{2}\sigma(a),$
from which the pairing * can be defined directly as
$h_{1}*h_{2}(m_{1}\otimes m_{2},m_{1}^{\prime}\otimes m_{2}^{\prime}):=\operatorname{Trd}_{A}\bigl{(}h_{1}(m_{1},m_{1}^{\prime}),\sigma(h_{2}(m_{2},m_{2}^{\prime}))\bigr{)}.$
We finish this section with a number of results for later use. We first consider [10, Proposition 4.9] in our context:
Lemma 3.11.
If $b,c\in\operatorname{Sym}(A,\sigma)$, then $\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma}\simeq\varphi_{b,c}$, where $(A,\varphi_{b,c})\in\mathfrak{H}(S,\iota)$ is given by
$\varphi_{b,c}\colon A\times A\to S,\ (x,y)\mapsto\operatorname{Trd}_{A}(\sigma(x)byc).$
Proof.
The form $\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma}$ is defined on the $S$-module $(A\otimes_{S}\prescript{\iota}{}{A})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A$. Since $A$ is a left $A\otimes_{S}\prescript{\iota}{}{A}$-module, the left action of $A\otimes_{S}\prescript{\iota}{}{A}$ induces an isomorphism of left $A\otimes_{S}\prescript{\iota}{}{A}$-modules (and thus of $S$-modules):
$f\colon(A\otimes_{S}\prescript{\iota}{}{A})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A\to A,\ x\otimes y\otimes a\mapsto(x\otimes y)\cdot_{\mathrm{ts}}a=xa\sigma(y).$
Using (3.8), we verify that it is the required isometry:
$\varphi_{b,c}($ $f(x_{1}\otimes y_{1}\otimes a_{1}),f(x_{2}\otimes y_{2}\otimes a_{2}))$
$=\varphi_{b,c}(x_{1}a_{1}\sigma(y_{1}),x_{2}a_{2}\sigma(y_{2}))=\operatorname{Trd}_{A}(y_{1}\sigma(a_{1})\sigma(x_{1})bx_{2}a_{2}\sigma(y_{2})c)$
$=\operatorname{Trd}_{A}(\sigma(a_{1})\sigma(x_{1})bx_{2}a_{2}\sigma(y_{2})cy_{1})=\operatorname{Trd}_{A}(\sigma(x_{1}a_{1})bx_{2}a_{2}\sigma(\sigma(y_{1})cy_{2}))$
$=\operatorname{Trd}_{A}(\langle b\rangle_{\sigma}(x_{1}a_{1},x_{2}a_{2})\sigma(\langle c\rangle_{\sigma}(y_{1},y_{2})))$
$=\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma}(x_{1}\otimes y_{1}\otimes a_{1},x_{2}\otimes y_{2}\otimes a_{2}).\qed$
Next we show that * is well-behaved under scalar extensions, and start with the following lemma (for which we could not find a reference):
Lemma 3.12.
Let $\Lambda$ be a commutative ring, and let $B$ and $C$ be $\Lambda$-algebras. Let $M_{1}$ be a right $B$-module and $M_{2}$ a left $B$-module. Then
$(M_{1}\otimes_{B}M_{2})\otimes_{\Lambda}C$ $\rightarrow(M_{1}\otimes_{\Lambda}C)\otimes_{B\otimes_{\Lambda}C}(M_{2}\otimes_{\Lambda}C),$
$(m_{1}\otimes m_{2})\otimes c$ $\mapsto(m_{1}\otimes 1)\otimes(m_{2}\otimes c)$
is an isomorphism of right $C$-modules.
Proof.
Let $f$ be the map defined in the statement of the lemma and let
$g\colon(M_{1}\otimes_{\Lambda}C)\otimes_{B\otimes_{\Lambda}C}(M_{2}\otimes_{\Lambda}C)$ $\rightarrow(M_{1}\otimes_{B}M_{2})\otimes_{\Lambda}C,$
$(m_{1}\otimes c_{1})\otimes(m_{2}\otimes c_{2})$ $\mapsto(m_{1}\otimes m_{2})\otimes c_{1}c_{2}.$
A standard (but lengthy) verification shows that $f$ and $g$ are well-defined and additive, are inverses of each other, and that $f$ is right $C$-linear. ∎
Lemma 3.13.
Let $T$ be a commutative $R$-algebra. Then
$(h_{1}*h_{2})\otimes_{R}T\simeq(h_{1}\otimes_{R}T)*(h_{2}\otimes_{R}T).$
Proof.
The form $(h_{1}*h_{2})\otimes_{R}T$ is defined on $((M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A)\otimes_{R}T$. Using Lemma 3.12 twice we have
$((M_{1}\otimes_{S}$ $\prescript{\iota}{}{M}_{2})\otimes_{A\otimes_{S}\prescript{\iota}{}{A}}A)\otimes_{R}T$
$\cong((M_{1}\otimes_{S}\prescript{\iota}{}{M}_{2})\otimes_{R}T)\otimes_{(A\otimes_{S}\prescript{\iota}{}{A})\otimes_{R}T}(A\otimes_{R}T)$
(3.10) $\cong((M_{1}\otimes_{R}T)\otimes_{S\otimes_{R}T}(\prescript{\iota}{}{M}_{2}\otimes_{R}T))\otimes_{(A\otimes_{R}T)\otimes_{S\otimes_{R}T}(\prescript{\iota}{}{A}\otimes_{R}T)}(A\otimes_{R}T),$
and the successive isomorphisms are given by
$[(m_{1}\otimes m_{2})\otimes a]\otimes t\mapsto[(m_{1}\otimes m_{2})\otimes 1]\otimes(a\otimes t)\mapsto[(m_{1}\otimes 1)\otimes(m_{2}\otimes 1)]\otimes(a\otimes t).$
We denote the composition of these two isomorphisms by $\xi$.
By definition, $(h_{1}*h_{2})\otimes_{R}T=(T_{\sigma}\bullet(h_{1}\otimes_{S}\prescript{\iota}{}{h}_{2}))\otimes_{R}T$, while
$(h_{1}\otimes_{R}T)*(h_{2}\otimes_{R}T)=T_{\sigma\otimes\operatorname{id}_{T}}\bullet((h_{1}\otimes_{R}T)\otimes_{S\otimes_{R}T}(\prescript{\iota}{}{h}_{2}\otimes_{R}T))$
and is defined on the module (3.10) (using that $Z(A\otimes_{R}T)=Z(A)\otimes_{R}T$, cf. Proposition 1.7). We check that $\xi$ is an isometry:
$T_{\sigma\otimes\operatorname{id}_{T}}\bullet((h_{1}\otimes_{R}T)\otimes_{S\otimes_{R}T}(\prescript{\iota}{}{h}_{2}\otimes_{R}T))(\xi(m_{1}\otimes m_{2}\otimes a\otimes t),\xi(m^{\prime}_{1}\otimes m^{\prime}_{2}\otimes a^{\prime}\otimes t^{\prime}))$
$=\operatorname{Trd}_{A\otimes T}\left((h_{1}\otimes T)((m_{1}\otimes 1)(a\otimes t),(m^{\prime}_{1}\otimes 1)(a^{\prime}\otimes t^{\prime}))\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.(\sigma\otimes\operatorname{id}_{T})((h_{2}\otimes T)(m_{2}\otimes 1,m^{\prime}_{2}\otimes 1))\right)$
$=\operatorname{Trd}_{A\otimes T}\left((h_{1}\otimes T)(m_{1}a\otimes t,m^{\prime}_{1}a^{\prime}\otimes t^{\prime})(\sigma\otimes\operatorname{id}_{T})((h_{2}\otimes T)(m_{2}\otimes 1,m^{\prime}_{2}\otimes 1))\right)$
$=\operatorname{Trd}_{A\otimes T}\left((h_{1}(m_{1}a,m^{\prime}_{1}a^{\prime})\otimes tt^{\prime})(\sigma\otimes\operatorname{id}_{T})(h_{2}(m_{2},m^{\prime}_{2})\otimes 1)\right)$
$=\operatorname{Trd}_{A\otimes T}(h_{1}(m_{1}a,m^{\prime}_{1}a^{\prime})\sigma(h_{2}(m_{2},m^{\prime}_{2}))\otimes tt^{\prime})$
$=\operatorname{Trd}_{A}(h_{1}(m_{1}a,m^{\prime}_{1}a^{\prime})\sigma(h_{2}(m_{2},m^{\prime}_{2})))\otimes tt^{\prime}$
$=((T_{\sigma}\bullet(h_{1}\otimes\prescript{\iota}{}{h}_{2}))\otimes_{R}T)((m_{1}\otimes m_{2}\otimes a)\otimes t,(m^{\prime}_{1}\otimes m^{\prime}_{2}\otimes a^{\prime})\otimes t^{\prime}),$
where we used that $\operatorname{Trd}_{A}$ commutes with scalar extension in the penultimate step. ∎
Remark 3.14.
Recall that when $S$ is not connected, the Witt groups $W^{\varepsilon}(A,\sigma)$ and $W^{\varepsilon}(S,\iota)$ are trivial, cf. Remark 1.16. Therefore the construction of the pairing * is not interesting in this case.
4. Pairings and PSD quadratic forms
Throughout this section we assume that the commutative ring $R$ (with $2\in R^{\times}$) is semilocal. Let $(A,\sigma)$ be an Azumaya algebra with involution over $R$, $S=Z(A)$ and $\iota=\sigma|_{S}$.
Let $(M,h)$ be a hermitian form over $(S,\iota)$. Then $h(x,x)\in\operatorname{Sym}(S,\iota)=R$ for all $x\in M$. Therefore we say that $h$ is positive semidefinite (resp. negative semidefinite) at $\alpha\in\operatorname{Sper}R$ if $h(x,x)\in\alpha$ (resp. $h(x,x)\in-\alpha$ ) for all $x\in M$. We use the standard abbreviations PSD and NSD.
Lemma 4.1.
Let $(M,h)$ be a hermitian form over $(S,\iota)$ and let $\alpha\in\operatorname{Sper}R$. Then $h$ is PSD at $\alpha$ if and only if $h\otimes\kappa(\alpha)$ is PSD at the ordering $\bar{\alpha}$ induced by $\alpha$ on $\kappa(\alpha)$.
Proof.
We use the description of $M\otimes_{R}\kappa(\alpha)$ given in Remark 1.17.
$\Rightarrow$”: Let $m\otimes\frac{1}{\bar{b}}\in M\otimes_{R}\kappa(\alpha)$, where $m\in M$ and $b\in R\setminus\operatorname{Supp}(\alpha)$. Then
$h\otimes\kappa(\alpha)(m\otimes\dfrac{1}{\bar{b}},m\otimes\dfrac{1}{\bar{b}})=\overline{h(m,m)}(\dfrac{1}{\bar{b}})^{2},$
which is in $\bar{\alpha}$ by hypothesis.
$\Leftarrow$”: Let $m\in M$. We have $h\otimes\kappa(\alpha)(m\otimes 1,m\otimes 1)=\overline{h(m,m)}$, which belongs to $\bar{\alpha}$, and thus $h(m,m)\in\alpha$. ∎
Remark 4.2.
Let $F$ be a real closed field and let $P$ be the unique ordering on $F$. Recall from Remark 2.6 and the beginning of Section 2 that if $(A,\sigma)=(M_{n}(D),\vartheta^{t})$ with $(D,\vartheta)\in\{(F,\operatorname{id}),(F(\sqrt{-1}),\gamma),((-1,-1)_{F},\gamma)\}$, where $\gamma$ denotes conjugation, resp. quaternion conjugation, and $a\in\operatorname{Sym}(A,\sigma)$, then $\operatorname{sign}^{\mathscr{M}_{P}}_{P}\langle a\rangle_{\sigma}$ is plus or minus the standard Sylvester signature of $a$ at $P$.
Indeed, if $a$ is a matrix in $\operatorname{Sym}(M_{n}(D),\vartheta^{t})$ such that $\vartheta^{t}(a)=a$, then under the hermitian Morita equivalence between $\mathfrak{Herm}^{\varepsilon}(M_{n}(D),\vartheta^{t})$ and $\mathfrak{Herm}^{\varepsilon}(D,\vartheta)$ in [4, equation (2.1)], the one-dimensional hermitian form $\langle a\rangle_{\vartheta^{t}}$ over $(M_{n}(D),\vartheta^{t})$ corresponds to the hermitian form over $(D,\vartheta)$ whose Gram matrix is $a$, and the signature of $\langle a\rangle_{\vartheta^{t}}$ is defined to be the signature of the matrix $a$. The definition of signature of hermitian forms allows the use of a different Morita equivalence $\mathscr{M}_{P}$ than this particular one, which may result in a change of sign.
In particular, there exists $\varepsilon\in\{-1,1\}$ (which depends only on $\mathscr{M}_{P}$) such that for all $a\in\operatorname{Sym}(M_{n}(D),\vartheta^{t})$, if $\operatorname{sign}^{\mathscr{M}_{P}}_{P}\langle a\rangle_{\vartheta^{t}}$ is maximal (among signatures of hermitian forms of the form $\langle b\rangle_{\vartheta^{t}}$), then $\varepsilon a$ is a PSD matrix.
Lemma 4.3.
Let $F$ be a field and $(B,\tau)$ a central simple $F$-algebra with involution. Let $P\in X_{F}\setminus\operatorname{Nil}[B,\tau]$. There is $\delta\in\{-1,1\}$ such that for every $b,c\in M^{\mathscr{M}_{P}}_{P}(B,\tau)$ and every $x\in B$, $\operatorname{Trd}_{B}(\tau(x)bxc)\in\delta\cdot P$.
Proof.
By [4, Proposition 6.7 and equation (6.1)], we have
$m_{P}(B,\tau)=m_{P^{\prime}}(B\otimes_{F}F_{P},\tau\otimes\operatorname{id}),$
where $P^{\prime}$ denotes the unique ordering on $F_{P}$. Hence, $x\in M^{\mathscr{M}_{P}}_{P}(B,\tau)$ implies $x\otimes 1\in M^{\mathscr{M}_{P^{\prime}}}_{P^{\prime}}(B\otimes_{F}F_{P},\tau\otimes\operatorname{id})$. Therefore, we can assume that $F$ is real closed with unique ordering $P$, that $(B,\tau)=(M_{n}(D),\operatorname{Int}(a)\circ\vartheta^{t})$ for some $a\in M_{n}(D)^{\times}$ with $\vartheta$ of the same kind as $\tau$ and $(D,\vartheta)\in\{(F,\operatorname{id}),(F(\sqrt{-1}),\gamma),((-1,-1)_{F},\gamma)\}$, where $\gamma$ denotes conjugation, resp. quaternion conjugation (cf. [3, First page of Section 2.3]). Note that $\vartheta(a)^{t}=\delta a$ for some $\delta\in\{-1,1\}$ by [17, Propositions 2.7 and 2.18]. Furthermore, the algebra $(B,\tau)$ is formally real in the sense of [4, Definition 3.2] since $P\not\in\operatorname{Nil}[B,\tau]$ (apply [4, Proposition 6.6], where $\widetilde{X}_{F}:=X_{F}\setminus\operatorname{Nil}[B,\tau]$), and it follows that we can choose the involution $\vartheta$ such that $\delta=1$ by [4, Corollary 3.8].
With reference to [1, Remark 3.13], observe that
$a^{-1}\langle b\rangle_{\tau}=\langle a^{-1}b\rangle_{\operatorname{Int}(a^{-1})\circ\tau}=\langle a^{-1}b\rangle_{\vartheta^{t}},$
and that there is a Morita equivalence $\mathscr{M}^{\prime}_{P}$ such that
$m_{P}(B,\tau)=\operatorname{sign}_{P}^{\mathscr{M}_{P}}\langle b\rangle_{\tau}=\operatorname{sign}_{P}^{\mathscr{M}^{\prime}_{P}}a^{-1}\langle b\rangle_{\tau}=\operatorname{sign}_{P}^{\mathscr{M}^{\prime}_{P}}\langle a^{-1}b\rangle_{\vartheta^{t}}.$
Similarly we have $\operatorname{sign}^{\mathscr{M}^{\prime}_{P}}_{P}\langle a^{-1}c\rangle_{\vartheta^{t}}=m_{P}(B,\tau)$. By Remark 4.2 there exists $\varepsilon\in\{-1,1\}$ such that $\varepsilon a^{-1}b$ and $\varepsilon a^{-1}c$ are both PSD with respect to $P$. Write $b^{\prime}:=\varepsilon a^{-1}b$ and $c^{\prime}:=\varepsilon a^{-1}c$. The matrix $\vartheta(a)^{t}c^{\prime}a$ is then PSD over $(D,\vartheta)$, and thus is a hermitian square in $(M_{n}(D),\vartheta^{t})$, i.e., there exists a matrix $c^{\prime\prime}\in M_{n}(D)$ such that $\vartheta(a)^{t}c^{\prime}a=\vartheta(c^{\prime\prime})^{t}c^{\prime\prime}$ (this is a classical consequence of the principal axis theorem, which also holds for quaternion matrices by [29, Corollary 6.2], cf. [4, Appendix A]). It follows that
$\operatorname{Trd}_{B}(\tau(x)bxc)$ $=\operatorname{Trd}_{B}((\operatorname{Int}(a)\circ\vartheta^{t})(x)(a\varepsilon b^{\prime})x(a\varepsilon c^{\prime}))=\operatorname{Trd}_{B}(a\vartheta(x)^{t}a^{-1}ab^{\prime}xac^{\prime})$
$=\operatorname{Trd}_{B}(a\vartheta(x)^{t}b^{\prime}xac^{\prime})=\delta\operatorname{Trd}_{B}(\vartheta(x)^{t}b^{\prime}x\vartheta(a)^{t}c^{\prime}a)$
$=\delta\operatorname{Trd}_{B}(\vartheta(x)^{t}b^{\prime}x\vartheta(c^{\prime\prime})^{t}c^{\prime\prime})=\delta\operatorname{Trd}_{B}(c^{\prime\prime}\vartheta(x)^{t}b^{\prime}x\vartheta(c^{\prime\prime})^{t}),$
which belongs to $\delta P$ since $b^{\prime}$ is PSD. ∎
Proposition 4.4.
Let $\alpha\in\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]$. Then there is $\delta\in\{-1,1\}$ such that for every $b,c\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$, the hermitian form $\delta(\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma})$ is nonsingular and PSD with respect to $\alpha$.
Proof.
The forms $\langle b\rangle_{\sigma}$ and $\langle c\rangle_{\sigma}$ are nonsingular since $b$ and $c$ are invertible, and thus $\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma}$ is nonsingular by Corollary 3.8. By Lemma 4.1, it suffices to show that the form $(\delta\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes\kappa(\alpha)$ is PSD with respect to $\bar{\alpha}$. Using now that $(\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes\kappa(\alpha)\simeq\langle b\otimes 1\rangle_{\sigma\otimes\operatorname{id}}*\langle c\otimes 1\rangle_{\sigma\otimes\operatorname{id}}$ by Lemma 3.13, and that $b\otimes 1,c\otimes 1\in M^{\mathscr{M}_{\bar{\alpha}}}_{\bar{\alpha}}(A(\alpha),\sigma(\alpha))$ by Corollary 2.19, we can assume that $(A,\sigma)$ is a central simple algebra with involution over the field $\kappa(\alpha)$ with ordering $\bar{\alpha}$ that is non-nil by Lemma 2.7. The result then follows from Lemma 4.3 since $\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma}$ is isometric to the hermitian form $\varphi_{b,c}$, cf. Lemma 3.11. ∎
Lemma 4.5.
Let $\alpha\in\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]$, and let $b,c\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$. Then there are nonsingular hermitian forms $\varphi_{1}$ and $\varphi_{2}$ over $(S,\iota)$ that are PSD at $\alpha$ such that $\varphi_{1}\otimes_{S}\langle b\rangle_{\sigma}\simeq\varphi_{2}\otimes_{S}\langle c\rangle_{\sigma}$.
Proof.
Let $\delta\in\{-1,1\}$ be as given by Proposition 4.4. Then the hermitian forms $\delta(\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma})$ and $\delta(\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma})$ over $(S,\iota)$ are PSD with respect to $\alpha$. Furthermore, they are nonsingular since $b$ and $c$ are invertible and by Corollary 3.8. By Theorem 3.9 we have
$(\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes_{S}\langle b\rangle_{\sigma}\simeq(\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes_{S}\langle c\rangle_{\sigma},$
hence
$\delta(\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes_{S}\langle b\rangle_{\sigma}\simeq\delta(\langle b\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes_{S}\langle c\rangle_{\sigma},$
and we conclude with Proposition 4.4. ∎
5. Sylvester’s law of inertia and Pfister’s local-global principle
Let $(A,\sigma)$ be an Azumaya algebra with involution over $R$, $S=Z(A)$ and $\iota=\sigma|_{S}$. In the case of central simple algebras with involution over fields, a weaker version of the following result appears in [4, Theorem 8.9].
Theorem 5.1 (Sylvester’s law of inertia).
Assume that $R$ is semilocal connected and let $\alpha\in\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]$. Then $S$ is connected and, letting $t:=\operatorname{rank}_{S}A$, we have:
  1. (1)
    Let $h$ be a nonsingular hermitian form over $(A,\sigma)$. For every $c\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$ there are $w_{1},\ldots,w_{t},u_{1},\ldots,u_{r},v_{1},\ldots,v_{s}\in\alpha\cap R^{\times}$ such that
    $\langle w_{1},\ldots,w_{t}\rangle\otimes_{R}h\simeq\langle u_{1},\ldots,u_{r}\rangle\otimes_{R}\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{s}\rangle\otimes_{R}\langle c\rangle_{\sigma}.$
  2. (2)
    Assume $\langle a_{1},\ldots,a_{r}\rangle_{\sigma}\perp\langle-b_{1},\ldots,-b_{s}\rangle_{\sigma}\simeq\langle a^{\prime}_{1},\ldots,a^{\prime}_{p}\rangle_{\sigma}\perp\langle-b^{\prime}_{1},\ldots,-b^{\prime}_{q}\rangle_{\sigma}$ with $a_{1},\ldots,a_{r}$, $b_{1},\ldots,b_{s},a^{\prime}_{1},\ldots,a^{\prime}_{p},b^{\prime}_{1},\ldots,b^{\prime}_{q}\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$. Then $r=p$ and $s=q$.
Proof.
We first observe that $S$ is connected by Remark 2.9. By Lemma 1.14 every nonsingular hermitian form over $(S,\iota)$ is diagonalizable. By Proposition 1.3, $A$ is Azumaya over $S$, hence a projective $S$-module and so $\operatorname{rank}_{S}A$ is defined, and is constant since $S$ is connected, cf. [26, p. 12].
(1) By Theorem 3.9 we have $(\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes_{S}h\simeq(h*\langle c\rangle_{\sigma})\otimes_{S}\langle c\rangle_{\sigma}$. Since $h*\langle c\rangle_{\sigma}$ is nonsingular by Corollary 3.8 and diagonalizable, we can write
$h*\langle c\rangle_{\sigma}\simeq\langle u_{1},\ldots,u_{r}\rangle_{\iota}\perp\langle-v_{1},\ldots,-v_{s}\rangle_{\iota},$
with $u_{1},\ldots,u_{r},v_{1},\ldots,v_{s}\in\alpha\cap R^{\times}$. Therefore
$(\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma})\otimes_{S}h$ $\simeq(h*\langle c\rangle_{\sigma})\otimes_{S}\langle c\rangle_{\sigma}$
$\simeq\langle u_{1},\ldots,u_{r}\rangle_{\iota}\otimes_{S}\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{s}\rangle_{\iota}\otimes_{S}\langle c\rangle_{\sigma}.$
By Proposition 4.4, the nonsingular hermitian form $\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma}$ over $(S,\iota)$ is PSD or NSD with respect to $\alpha$. Up to replacing it by its opposite, we can assume it is PSD with respect to $\alpha$. As observed above, it is diagonalizable, and by Lemma 3.11 it is defined on $A$ and therefore has dimension $\operatorname{rank}_{S}A=t$. Hence there are $w_{1},\ldots,w_{t}\in\alpha\cap R^{\times}$ such that $\langle c\rangle_{\sigma}*\langle c\rangle_{\sigma}\simeq\langle w_{1},\ldots,w_{t}\rangle_{\iota}$, and thus
$\langle w_{1},\ldots,w_{t}\rangle_{\iota}\otimes_{S}h\simeq\langle u_{1},\ldots,u_{r}\rangle_{\iota}\otimes_{S}\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{s}\rangle_{\iota}\otimes_{S}\langle c\rangle_{\sigma}.$
The result now follows from Lemma 1.1.
(2) For dimension reasons (after localization, since $A$ is a projective $R$-module) we have $r+s=p+q$. The result will follow if we show $r-s=p-q$, i.e., $r+q=p+s$. We have the following equality in the Witt group $W(A,\sigma)$:
$\langle a_{1},\ldots,a_{r},b^{\prime}_{1},\ldots,b^{\prime}_{q}\rangle_{\sigma}=\langle a^{\prime}_{1},\ldots,a^{\prime}_{p},b_{1},\ldots,b_{s}\rangle_{\sigma},$
which implies that
$\langle a_{1},\ldots,a_{r},b^{\prime}_{1},\ldots,b^{\prime}_{q}\rangle_{\sigma}\perp\varphi\simeq\langle a^{\prime}_{1},\ldots,a^{\prime}_{p},b_{1},\ldots,b_{s}\rangle_{\sigma}\perp\psi,$
where $\varphi$ and $\psi$ are hyperbolic forms over $(A,\sigma)$. Taking signatures on both sides yields $(r+q)m_{\alpha}(A,\sigma)=(p+s)m_{\alpha}(A,\sigma)$ since hyperbolic forms have signature zero. The result follows. ∎
For $r_{1},\ldots,r_{\ell}\in R$, we use the notation $\langle\!\langle r_{1},\ldots,r_{\ell}\rangle\!\rangle$ to denote the Pfister form $\langle 1,r_{1}\rangle\otimes_{R}\cdots\otimes_{R}\langle 1,r_{\ell}\rangle$.
Corollary 5.2.
Assume that $R$ is semilocal connected.
  1. (1)
    Assume that $\sigma$ is of orthogonal or symplectic type (so that $(S,\iota)=(R,\operatorname{id})$). Let $\alpha\in\operatorname{Nil}[A,\sigma]$ and let $h$ be a nonsingular hermitian form over $(A,\sigma)$. Then there is a nonsingular quadratic form $q$ over $R$ of dimension $\operatorname{rank}_{R}A$ that is PSD at $\alpha$ and such that $q\otimes_{R}h$ is hyperbolic.
  2. (2)
    Let $\alpha\in\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]$, let $a,b\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$, and let $t:=\operatorname{rank}_{S}A$. Then there are $\ell\in\mathbb{N}$ and $r_{1},\ldots,r_{\ell},w_{1},\ldots,w_{t}\in\alpha\cap R^{\times}$ such that
    $(\langle w_{1},\ldots,w_{t}\rangle\otimes_{R}\langle\!\langle r_{1},\ldots,r_{\ell}\rangle\!\rangle)\otimes_{R}\langle a,-b\rangle_{\sigma}$
    is hyperbolic.
Proof.
(1) By Lemma 2.11 there exists a skew-symmetric element $a\in A^{\times}$. Let $\tau=\operatorname{Int}(a)\circ\sigma$ and note that $\tau$ is orthogonal if $\sigma$ is symplectic and vice versa (indeed, by Proposition 1.12 it suffices to check it for the central simple $\kappa(\alpha)$-algebra with involution $(A(\alpha),\sigma(\alpha))$, where it is true by [17, Proposition 2.7]). It follows from Remark 2.6 and Lemma 2.7 (1$\Leftrightarrow$(2) that $\alpha\in\operatorname{Sper}(R)\setminus\operatorname{Nil}[A,\tau]$. Scaling $h$ by $a$ gives the nonsingular skew-hermitian form $h^{\prime}:=ah$ over $(A,\tau)$.
Let $c\in M^{\mathscr{M}}_{\alpha}(A,\tau)$. Then the form $\langle c\rangle_{\tau}$ is nonsingular, and so the pairing $\langle c\rangle_{\tau}*\langle c\rangle_{\tau}$ is also nonsingular by Corollary 3.8. By Proposition 4.4, there exists $\delta\in\{-1,1\}$ such that the form $\delta(\langle c\rangle_{\tau}*\langle c\rangle_{\tau})$ is PSD at $\alpha$.
We now consider the form $\delta(\langle c\rangle_{\tau}*\langle c\rangle_{\tau})\otimes_{R}h^{\prime}$ which is isometric to $\delta(h^{\prime}*\langle c\rangle_{\tau})\otimes_{R}\langle c\rangle_{\tau}$ by Theorem 3.9. But $h^{\prime}*\langle c\rangle_{\tau}$ is a pairing of a skew-hermitian and a hermitian form over $(A,\tau)$, and so is skew-symmetric over $(S,\iota)=(R,\operatorname{id})$ by (3.7). Hence, $h^{\prime}*\langle c\rangle_{\tau}$ is hyperbolic by [16, I, Corollary 4.1.2] since every projective $R$-module is free by [11].
Therefore, letting $q:=\delta(\langle c\rangle_{\tau}*\langle c\rangle_{\tau})$, the form $q\otimes_{R}h^{\prime}$ is hyperbolic and since scaling by $a^{-1}$ commutes with tensoring by $q$, we obtain that $q\otimes_{R}h$ is hyperbolic by [16, I, Theorem 9.3.5] applied to the scaling-by-$a^{-1}$ Morita equivalence. Note that $q$ is diagonalizable since $R$ is semilocal and $A$ is projective over $R$, and that the dimension of $q$ is $\operatorname{rank}_{R}A$, cf. Lemma 3.11.
(2) Observe that $\operatorname{rank}_{S}A$ is constant by the first paragraph of the proof of Theorem 5.1. Let $c\in M^{\mathscr{M}}_{\alpha}(A,\sigma)$. By Theorem 5.1, there are $w_{1},\ldots,w_{t},u_{1},\ldots,u_{r},v_{1},\ldots,v_{s}\in\alpha\cap R^{\times}$ such that
$\langle w_{1},\ldots,w_{t}\rangle\otimes_{R}\langle a,-b\rangle_{\sigma}\simeq\langle u_{1},\ldots,u_{r}\rangle\otimes_{R}\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{s}\rangle\otimes_{R}\langle c\rangle_{\sigma}.$
The signature at $\alpha$ of the left-hand side is 00, so we must have $r=s$. Therefore,
$\langle w_{1},\ldots,w_{t}\rangle\otimes_{R}\langle a,-b\rangle_{\sigma}$ $\simeq\langle u_{1},\ldots,u_{r}\rangle\otimes_{R}\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{r}\rangle\otimes_{R}\langle c\rangle_{\sigma}$
$\simeq\langle\bar{u}\rangle\otimes_{R}\langle c\rangle_{\sigma}\perp\langle-\bar{v}\rangle\otimes_{R}\langle c\rangle_{\sigma}\simeq\langle\bar{u},-\bar{v}\rangle\otimes_{R}\langle c\rangle_{\sigma},$
where we write $\bar{u}$ for $u_{1},\ldots,u_{r}$ and similarly for $\bar{v}$. We use the notation $\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle$ for the Pfister form $\langle\!\langle u_{1},\ldots,u_{r},v_{1},\ldots,v_{s}\rangle\!\rangle$. The form $\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle\otimes_{R}\langle\bar{u},-\bar{v}\rangle$ is hyperbolic since
$\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle\otimes_{R}\langle\bar{u},-\bar{v}\rangle$ $\simeq\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle\otimes_{R}\langle\bar{u}\rangle\perp\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle\otimes_{R}\langle-\bar{v}\rangle$
[…]
$\simeq r\times\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle\perp-r\times\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle,$
where the final equality holds since Pfister forms over $R$ are round (cf. [5, Corollary 2.16] which uses the fact that the hypothesis $2\in R^{\times}$ ensures that $|R/\mathfrak{m}|>2$ for every maximal ideal $\mathfrak{m}$ of $R$). Therefore $(\langle w_{1},\ldots,w_{t}\rangle\otimes_{R}\langle\!\langle\bar{u},\bar{v}\rangle\!\rangle)\otimes_{R}\langle a,-b\rangle_{\sigma}$ is hyperbolic, proving the result. ∎
The main idea of the proof of the next result comes from Marshall’s proof of Pfister’s local-global principle in [23, Theorem 4.12].
Proposition 5.3.
Let $R$ be semilocal connected and let $t_{0}:=\operatorname{rank}_{R}A$. Let $h$ be a nonsingular hermitian form over $(A,\sigma)$ such that for every $n\in\mathbb{N}$, the form $2^{n}t_{0}^{2}\times h$ is not hyperbolic. Then there is $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$ such that $\operatorname{sign}^{\mathscr{M}}_{\alpha}h\not=0$.
Proof.
Observe that if $t=\operatorname{rank}_{S}A$ is defined, then $t_{0}$ is equal to $t$ (if $R=S$) or $2t$ (if $R\not=S$). We identify nonsingular hermitian forms with their classes in the Witt group $W(A,\sigma)$, considered as a $W(R)$-module. Note that a nonsingular hermitian form is hyperbolic if and only if its class is zero since we assume that $2\in A^{\times}$ (and therefore metabolic forms are hyperbolic).
For the convenience of the reader we first present the ideas of the three main steps of the proof before giving the full details:
  1. (1)
    We define a maximal non-empty set $\mathcal{S}$ of (nonsingular) Pfister forms over $R$ such that $p\cdot h\not=0$ for every $p\in\mathcal{S}$, and such that $\mathcal{S}$ is closed under products. This final property produces an ideal $J:=\bigcup_{p\in\mathcal{S}}\operatorname{Ann}_{W(A,\sigma)}(p)$, and $h\not\in J$ by construction (this is linked to the notion of Pfister quotient in [23, Chapter 4.7]).
    The actual construction in the proof below is slighly different (a factor $t_{0}^{2}$ appears for technical reasons), and the ideal is called $J_{\mathcal{S}}$.
  2. (2)
    The maximality of $\mathcal{S}$ ensures that the set $\alpha_{0}$ of all elements of $R^{\times}$ represented by these Pfister forms is “almost” an ordering, more precisely: $\alpha_{0}=\alpha\cap R^{\times}$ for some $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$.
  3. (3)
    The final step of the proof consists in checking that $\operatorname{sign}^{\mathscr{M}}_{\alpha}\subseteq J$, proving the result since $h\not\in J$.
We now proceed with the proof. Let $\mathcal{S}\subseteq W(R)$ be a maximal subset of Pfister forms such that:
  • (a)
    $\mathcal{S}\cdot\mathcal{S}\subseteq\mathcal{S}$;
  • (b)
    for every $n\in\mathbb{N}$, $2^{n}\times\langle 1\rangle\in\mathcal{S}$;
  • (c)
    for every $p\in\mathcal{S}$, $(t_{0}^{2}\times p)\cdot h\not=0$.
Observe that $\mathcal{S}$ exists since the set $\{2^{n}\times\langle 1\rangle\mid n\in\mathbb{N}\}$ satisfies all these conditions. Define
$\alpha_{0}:=\{r\in R^{\times}\mid\exists p\in\mathcal{S}\quad\langle r\rangle\cdot p=p\}.$
Clearly $\alpha_{0}\cdot\alpha_{0}\subseteq\alpha_{0}$. We first prove the following four properties:
  1. (P1)
    $\alpha_{0}\cup-\alpha_{0}=R^{\times}$;
  2. (P2)
    $\alpha_{0}\cap-\alpha_{0}=\emptyset$;
  3. (P3)
    for every $u\in\alpha_{0}$, $\langle 1,u\rangle\in\mathcal{S}$;
  4. (P4)
    for every $k\in\mathbb{N}$ and all $u_{1},\ldots,u_{k}\in\alpha_{0}$, $\langle\!\langle u_{1},\ldots,u_{k}\rangle\!\rangle\in\mathcal{S}$.
Proof of (P1): For $x\in R^{\times}$ we define $\mathcal{S}_{x}:=\mathcal{S}\cup\langle 1,x\rangle\mathcal{S}$. Obviously $\mathcal{S}\subseteq\mathcal{S}_{x}$. In particular $\mathcal{S}_{x}$ satisfies property (b). We check the non-obvious case of property (a): If $p,q\in\mathcal{S}$, then $\langle 1,x\rangle p\langle 1,x\rangle q=2\langle 1,x\rangle pq$, which belongs to $\mathcal{S}_{x}$ by properties (a) and (b) of $\mathcal{S}$.
Let $x\in R^{\times}$. If $x\not\in\alpha_{0}$ then $\langle 1,x\rangle\not\in\mathcal{S}$ (otherwise, since $\langle x\rangle\langle 1,x\rangle=\langle 1,x\rangle$ we would get $x\in\alpha_{0}$), so $\mathcal{S}\subsetneq\mathcal{S}_{x}$.
So if we assume that $x\not\in\alpha_{0}$ and $-x\not\in\alpha_{0}$ we obtain $\mathcal{S}\subsetneq\mathcal{S}_{x}$ and $\mathcal{S}\subsetneq\mathcal{S}_{-x}$. Since both $\mathcal{S}_{x}$ and $\mathcal{S}_{-x}$ satisfy properties (a) and (b), by maximality of $\mathcal{S}$ we get that (c) does not hold for either of $\mathcal{S}_{x}$ and $\mathcal{S}_{-x}$: There are $p,q\in\mathcal{S}$ such that $t_{0}^{2}\times\langle 1,x\rangle ph=0$ and $t_{0}^{2}\times\langle 1,-x\rangle qh=0$. Therefore $t_{0}^{2}\times\langle 1,x\rangle pqh=0$ and $t_{0}^{2}\times\langle 1,-x\rangle pqh=0$. Adding both we obtain $t_{0}^{2}\times 2pqh=0$ in $W(A,\sigma)$, a contradiction since $2pq\in\mathcal{S}$. End of the proof of (P1).
Proof of (P2): Assume that there is $x\in\alpha_{0}\cap-\alpha_{0}$. Then there are $p,q\in\mathcal{S}$ such that $\langle x\rangle p=p$ and $\langle-x\rangle q=q$. Therefore $\langle x\rangle pq=pq$ and $\langle-x\rangle pq=pq$. Adding both, we get $0=2pq$ and thus $02pq\cdot h=$ 0, a contradiction since $2pq\in\mathcal{S}$. End of the proof of (P2).
Proof of (P3): Let $u\in\alpha_{0}$ and let $p_{0}\in\mathcal{S}$ be such that $\langle u\rangle\cdot p_{0}=p_{0}$. Consider, as in the proof of (P1), $\mathcal{S}_{u}:=\mathcal{S}\cup\langle 1,u\rangle\mathcal{S}$. As seen in the proof of (P1), $\mathcal{S}_{u}$ satisfies properties (a) and (b). We check property (c): Assume $t_{0}^{2}\times\langle 1,u\rangle ph=0$ for some $p\in\mathcal{S}$. Then $t_{0}^{2}\times\langle 1,u\rangle p_{0}ph=0$, i.e., $t_{0}^{2}\times 2p_{0}ph=0$, which is impossible since $2p_{0}p\in\mathcal{S}$. Since $\mathcal{S}_{u}$ satisfies properties (a), (b) and (c), and contains $\mathcal{S}$, we must have $\mathcal{S}=\mathcal{S}_{u}$ by maximality of $\mathcal{S}$. Therefore $\langle 1,u\rangle\in\mathcal{S}$. End of the proof of (P3).
(P4) is a direct consequence of (P3) and property (a). This finishes the proof of (P1)–(P4).
Next, we define
$J_{\mathcal{S}}:=\{\psi\in W(A,\sigma)\mid\exists p\in\mathcal{S}\quad(t_{0}^{2}\times p)\cdot\psi=0\}.$
Clearly, $h\not\in J_{\mathcal{S}}$ by property (c).
Consider the map
$\chi\colon R^{\times}\rightarrow\{-1,1\},\quad\chi(r):=\begin{cases}1&\text{if}r\in\alpha_{0}\\ -1&\text{if }r\in-\alpha_{0}\end{cases}.$
Then $\chi$ is a signature of $R$. Indeed, it is a character on $R^{\times}$ since $\alpha_{0}\cdot\alpha_{0}\subseteq\alpha_{0}$ and by properties (P1) and (P2). Furthermore, by [14, top of p. 88], $\chi$ will then be a signature of $R$ if it satisfies $\chi(-1)=-1$ (which is true), as well as $\chi(r_{1}^{2}f_{1}+\cdots+r_{k}^{2}f_{k})=1$ whenever $f_{1},\ldots,f_{k}\in R^{\times}$ with $\chi(f_{1})=\cdots=\chi(f_{k})=1$ and $r_{1},\ldots,r_{k}\in R$ are such that $z:=r_{1}^{2}f_{1}+\cdots+r_{k}^{2}f_{k}\in R^{\times}$. We check this: Since $f_{1},\ldots,f_{k}\in\alpha_{0}$, we have by (P4) that the Pfister form $\langle\!\langle f_{1},\ldots,f_{k}\rangle\!\rangle$ is in $\mathcal{S}$. Clearly, $z\in D_{R}\langle\!\langle f_{1},\ldots,f_{k}\rangle\!\rangle$ and by [5, p. 94, Theorem 2.1] (which applies since the hypothesis $2\in R^{\times}$ ensures that every quotient of $R$ by a proper ideal has at least 3 elements) we have $\langle z\rangle\langle\!\langle f_{1},\ldots,f_{k}\rangle\!\rangle=\langle\!\langle f_{1},\ldots,f_{k}\rangle\!\rangle$. Thus $z\in\alpha_{0}$ and $\chi(z)=1$.
By the correspondence between $\operatorname{Sign}R$ and $\operatorname{Sper}^{\mathrm{max}}R$ (cf. Section 1.4) there is $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$ such that $\chi=\operatorname{sign}_{\alpha}$, and thus $\alpha_{0}=\alpha\cap R^{\times}$.
Claim 1: If $\sigma$ is of unitary type, then $\alpha\not\in\operatorname{Nil}[A,\sigma]$.
Proof of Claim 1: Observe first that for every $q\in\mathcal{S}$, $q\otimes_{R}\langle 1\rangle_{\iota}$ is not hyperbolic (if it were, then $(q\otimes_{R}\langle 1\rangle_{\iota})\otimes_{S}h$ would be hyperbolic, and thus, by Lemma 1.1, $(q\otimes_{R}\langle 1\rangle_{\iota})\otimes_{S}h\simeq q\otimes_{R}h$ would also be hyperbolic, contradicting the definition of $\mathcal{S}$).
Assume that $\alpha\in\operatorname{Nil}[A,\sigma]=\operatorname{Nil}[S,\iota]=\mathring{H}(d)$ for some $d\in R^{\times}$ (by Lemma 2.7 and Corollary 2.8). Then $d\in\alpha\cap R^{\times}=\alpha_{0}$ and $\operatorname{sign}^{\mathscr{M}}\langle 1,d\rangle\otimes_{R}\langle 1\rangle_{\iota}=0$ on $\operatorname{Sper}R$. (Indeed, the signature is zero on $\operatorname{Nil}[A,\sigma]$ by definition, and the signature of $\langle 1,d\rangle$ is zero on $\operatorname{Sper}R\setminus\operatorname{Nil}[A,\sigma]$ by definition of $d$ and since $d\not\in\operatorname{Supp}\alpha$.) Note that $\langle 1,d\rangle\in\mathcal{S}$ by (P3). Since $R$ is connected, $S$ is quadratic étale over $R$ and $\iota$ is the standard involution on $S$ by Proposition 1.12, and we may use Lemma 2.4. Therefore, $\operatorname{sign}\operatorname{Tr}_{S/R}(\langle 1,d\rangle\otimes_{R}\langle 1\rangle_{\iota})=0$ on $\operatorname{Sper}R$. Since $\operatorname{Tr}_{S/R}(\langle 1,d\rangle\otimes_{R}\langle 1\rangle_{\iota})$ is nonsingular (cf. Section 1.2), Pfister’s local-global principle for semilocal rings (cf. [22, p. 194] or [5, Theorem 7.16]) applies and there is $k\in\mathbb{N}$ such that $2^{k}\times\operatorname{Tr}_{S/R}(\langle 1,d\rangle\otimes_{R}\langle 1\rangle_{\iota})$ is hyperbolic. Applying [8, Corollary 8.3] yields that $2^{k}\times\langle 1,d\rangle\otimes_{R}\langle 1\rangle_{\iota}$ is hyperbolic, a contradiction (as observed above) since $2^{k}\times\langle 1,d\rangle$ is in $\mathcal{S}$. End of the proof of Claim 1.
Claim 2: $\ker\operatorname{sign}^{\mathscr{M}}_{\alpha}\subseteq J_{\mathcal{S}}$.
Proof of Claim 2: Let $\psi\in\ker\operatorname{sign}^{\mathscr{M}}_{\alpha}$. We consider two cases.
  1. (1)
    $\alpha\in\operatorname{Nil}[A,\sigma]$. By Claim 1 we know that $\sigma$ must be of orthogonal or symplectic type. In particular $S=R$ and $t_{0}=t$. By Corollary 5.2(1), there is a nonsingular diagonal quadratic form $\langle u_{1},\ldots,u_{t_{0}}\rangle$ of dimension $t_{0}$ with coefficients in $\alpha$ (and thus in $\alpha\cap R^{\times}=\alpha_{0}$) such that $\langle u_{1},\ldots,u_{t_{0}}\rangle\psi$ is hyperbolic. Multiplying by $\langle\!\langle u_{1},\ldots,u_{t_{0}}\rangle\!\rangle$ and using that $\langle u_{i}\rangle\langle\!\langle u_{1},\ldots,u_{t_{0}}\rangle\!\rangle=\langle\!\langle u_{1},\ldots,u_{t_{0}}\rangle\!\rangle$ (cf. [5, Corollary 2.16] which uses the fact that the hypothesis $2\in R^{\times}$ ensures that $|R/\mathfrak{m}|>2$ for every maximal ideal $\mathfrak{m}$ of $R$), we obtain $t_{0}\times\langle\!\langle u_{1},\ldots,u_{t_{0}}\rangle\!\rangle\psi=0$. By property (P4) the form $\langle\!\langle u_{1},\ldots,u_{t_{0}}\rangle\!\rangle$ is in $\mathcal{S}$ and thus $\psi\in{J_{\mathcal{S}}}$.
  2. (2)
    $\alpha\not\in\operatorname{Nil}[A,\sigma]$. By Theorem 5.1, $t=\operatorname{rank}_{S}A$ exists and we have
    $\langle w_{1},\ldots,w_{t}\rangle\psi\simeq\langle u_{1},\ldots,u_{r}\rangle\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{s}\rangle\langle c\rangle_{\sigma},$
    for some $w_{1},\ldots,w_{t},u_{1},\ldots,u_{r},v_{1},\ldots,v_{s}\in\alpha\cap R^{\times}$ and $c\in\operatorname{Sym}(A^{\times},\sigma)$ such that $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle c\rangle_{\sigma}=m_{\alpha}(A,\sigma)$. Multiplying both sides by $\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle$ and then by $2$, the left-hand side becomes first $t\times\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle\psi$ (using that $\langle w_{i}\rangle\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle=\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle$) and then $t_{0}\times\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle\psi$, while the right-hand side still retains the same shape (up to taking larger values for $r$ and $s$, and different elements $u_{i},v_{j}\in\alpha\cap R^{\times}$). We thus have
    $t_{0}\times\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle\psi\simeq\langle u_{1},\ldots,u_{r}\rangle\langle c\rangle_{\sigma}\perp\langle-v_{1},\ldots,-v_{s}\rangle\langle c\rangle_{\sigma},$
    where $u_{1},\ldots,u_{r},v_{1},\ldots,v_{s}\in\alpha\cap R^{\times}$.
    Since $\operatorname{sign}^{\mathscr{M}}_{\alpha}\psi=0$ and $\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle u_{i}\rangle\langle c\rangle_{\sigma}=\operatorname{sign}^{\mathscr{M}}_{\alpha}\langle v_{j}\rangle\langle c\rangle_{\sigma}=m_{\alpha}(A,\sigma)$ for $i=1,\ldots,r$ and $j=1,\ldots,s$, we must have $r=s$. In particular we can pair each $u_{i}c$ with the corresponding $-v_{i}c$, so that
[…]
    Fact: For each $i=1,\ldots,r$ there is $p_{i}\in\mathcal{S}$ such that $t_{0}\times p_{i}\cdot\langle u_{i}c,-v_{i}c\rangle_{\sigma}=0$.
    Proof of the Fact: By Corollary 5.2(2), there are $z_{1},\ldots,z_{t}$, $r_{1},\ldots,r_{\ell}\in\alpha\cap R^{\times}=\alpha_{0}$ such that $\langle z_{1},\ldots,z_{t}\rangle\langle\!\langle r_{1},\ldots,r_{\ell}\rangle\!\rangle\langle u_{i}c,-v_{i}c\rangle_{\sigma}=0$. Multiplying by the Pfister form $\langle\!\langle z_{1},\ldots,z_{t}\rangle\!\rangle$, we obtain
    $(t\times\langle\!\langle z_{1},\ldots,z_{t},r_{1},\ldots,r_{\ell}\rangle\!\rangle)\langle u_{i}c,-v_{i}c\rangle_{\sigma}=0,$
    and thus $(t_{0}\times\langle\!\langle z_{1},\ldots,z_{t},r_{1},\ldots,r_{\ell}\rangle\!\rangle)\langle u_{i}c,-v_{i}c\rangle_{\sigma}=0$. The Fact follows since the form $\langle\!\langle z_{1},\ldots,z_{t},r_{1},\ldots,r_{\ell}\rangle\!\rangle$ is in $\mathcal{S}$ by property (P4). End of the proof of the Fact.
    Multiplying (5.1) by $t_{0}\times p_{1}\cdots p_{r}$ gives $t_{0}^{2}\times(p_{1}\cdots p_{r}\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle)\cdot\psi=0$, proving that $\psi$ is in $J_{\mathcal{S}}$ since $p_{1}\cdots p_{r}\langle\!\langle w_{1},\ldots,w_{t}\rangle\!\rangle\in\mathcal{S}$ (by properties (P4) and (a)). End of the proof of Claim 2.
The conclusion is now clear since $h\not\in J_{\mathcal{S}}$. ∎
Proposition 5.4.
Let $R$ be semilocal connected. The torsion in $W(A,\sigma)$ is $2$-primary.
Proof.
By [6, Theorem 8.7 and Remark 8.8] there exists a connected finite étale $R$-algebra $R_{1}$ of odd rank, and an Azumaya algebra with involution $(A_{1},\sigma_{1})$ over $R_{1}$ such that $A\otimes_{R}R_{1}$ and $A_{1}$ are Brauer equivalent over $S_{1}:=S\otimes_{R}R_{1}$, $\sigma$ and $\sigma_{1}$ are both unitary, or are both non-unitary, and such that at least one of the following holds:
  1. (1)
    $Z(A_{1})\cong R_{1}\times R_{1}$;
  2. (2)
    $\deg A_{1}=1$;
  3. (3)
    The index and degree of $A_{1}$ are equal and divide the index of $A$. Moreover, $\deg A_{1}$ is a power of $2$ and there exist $u,v\in A_{1}^{\times}$ such that $u^{2}\in R_{1}^{\times}$, $\sigma_{1}(u)=-u$, $\sigma_{1}(v)=-v$ and $uv=-vu$.
By [6, Corollary 7.4], the canonical map of Witt groups
$W(A,\sigma)\to W(A\otimes_{R}R_{1},\sigma\otimes\operatorname{id}_{R_{1}})$
is injective and thus it suffices to show that the torsion in $W(A\otimes_{R}R_{1},\sigma\otimes\operatorname{id}_{R_{1}})$ is $2$-primary.
If $S_{1}$ is not connected, then $W(A\otimes_{R}R_{1},\sigma\otimes\operatorname{id}_{R_{1}})=0$ (cf. Remark 1.16) and we can conclude. Thus we may assume that $S_{1}$ is connected, and in particular that we are not in case (1) above. By Theorem 1.15 we have a Witt group isomorphism
$W(A\otimes_{R}R_{1},\sigma\otimes\operatorname{id}_{R_{1}})\cong W^{\delta}(A_{1},\sigma_{1})$
for some $\delta\in\{-1,1\}$, where we may take $\delta=1$ if $\sigma$ and $\sigma_{1}$ are unitary, observing that $\sigma$ and $\sigma\otimes\operatorname{id}_{R_{1}}$ are of the same type. We now examine the remaining relevant cases from [6, Theorem 8.7], as listed in (2) and (3) above:
  1. (2)
    $\deg A_{1}=1$, i.e., $A_{1}=S_{1}$: Assume first that $\sigma_{1}$ is not unitary. By Proposition 1.12 we then have $A_{1}=S_{1}=R_{1}$ and $\sigma_{1}=\operatorname{id}_{S_{1}}$. In this case $W^{-1}(R_{1},\operatorname{id})=0$ by [16, I, Corollary 4.1.2] (whose hypotheses are satisfied since $R_{1}$ is connected and also semilocal by [16, VI, Proposition 1.1.1]), while the torsion in $W(R_{1},\operatorname{id})$ is $2$-primary by [5, Chapter V, Theorem 6.6].
    On the other hand, if $\sigma_{1}$ is unitary (so that we may take $\delta=1$), then $S_{1}$ is a quadratic étale $R_{1}$-algebra and $\sigma_{1}$ is the standard involution. By [8, Corollary 8.3], the map $h\mapsto\operatorname{Tr}_{S_{1}/R_{1}}\circ h$ is an injection from $W(S_{1},\sigma_{1})$ into $W(R_{1},\operatorname{id})$, which has $2$-primary torsion as observed above.
  2. (3)
    $\deg A_{1}$ is a power of $2$ and there exists $u\in A_{1}^{\times}$ such that $\sigma_{1}(u)=-u$. In particular, $\operatorname{rank}_{S_{1}}A_{1}$ and hence $\operatorname{rank}_{R_{1}}A_{1}$ are powers of $2$. We consider two cases:
    If $\delta=1$, we conclude with Proposition 5.3: If $h$ is torsion in $W(A_{1},\sigma_{1})$, then $h$ has zero signature at every ordering of $R_{1}$ and by Proposition 5.3, there is $n\in\mathbb{N}$ such that $2^{n}\times h$ is hyperbolic.
    If $\delta=-1$, we have
    $W^{-1}(A_{1},\sigma_{1})\cong W(A_{1},\operatorname{Int}(u)\circ\sigma_{1})$
    (by Morita equivalence, more precisely the $\mu$-conjugation equivalence of categories in [8, Section 2.7]), and we conclude again with Proposition 5.3 applied to $(A_{1},\operatorname{Int}(u)\circ\sigma_{1})$.∎
Remark 5.5.
If $R$ is semilocal with only $k$ maximal ideals, then any expression of $R$ as a product $R_{1}\times\cdots\times R_{n}$ of rings must be such that $n\leq k$. Therefore there is such an expression of $R$ as a product $R_{1}\times\cdots\times R_{t}$ that cannot be further decomposed as a product and thus where each $R_{i}$ is connected.
Theorem 5.6 (Pfister’s local-global principle).
Let $R$ be semilocal, and recall that we assume $2\in R^{\times}$. Let $(M,h)$ be a nonsingular hermitian form over $(A,\sigma)$. The following statements are equivalent:
  1. (1)
    $\operatorname{sign}^{\mathscr{M}}_{\alpha}h=0$ for every $\alpha\in\operatorname{Sper}R$.
  2. (2)
    $\operatorname{sign}^{\mathscr{M}}_{\alpha}h=0$ for every $\alpha\in\operatorname{Sper}^{\mathrm{max}}R$.
  3. (3)
    There exists $n\in\mathbb{N}\cup\{0\}$ such that $2^{n}\times h$ is hyperbolic.
In particular, the torsion in $W(A,\sigma)$ is $2$-primary.
Proof.
Observe that the final statement clearly follows from the equivalence of (1) and (3), since a torsion form has zero signature at every ordering.
Clearly (1) implies (2), and (3) implies (1), so we only need to show that (2) implies (3).
Following Remark 5.5, we may assume that $R=R_{1}\times\cdots\times R_{t}$ with $R_{1},\ldots,R_{t}$ connected semilocal rings. Writing $e_{1}=(1,0,\ldots,0),\ldots,e_{t}=(0,\ldots,0,1)$ in $R$, we have
$(A,\sigma)\cong(Ae_{1},\sigma|_{Ae_{1}})\times\cdots\times(Ae_{t},\sigma|_{Ae_{t}}).$
Furthermore, we can identify $M$ with $\bigoplus_{i=1}^{t}Me_{i}$, and we consider $h_{i}:=h|_{Me_{i}}$ as a hermitian form over $(Ae_{i},\sigma|_{Ae_{i}})$ for $i=1,\ldots,t$. A direct verification shows that:
  • $h$ is nonsingular if and only if each $h_{i}$ is nonsingular for $i=1,\ldots,t$ (using for instance that $h$ is nonsingular if and only if for every maximal ideal $\mathfrak{m}$ of $R$ the form $h\otimes_{R}R/{\mathfrak{m}}$ is nonsingular, cf. [16, I, Lemma 7.1.3]).
  • If $h_{i}$ is hyperbolic for $i=1,\ldots,t$, then $h$ is hyperbolic. Indeed: We can write $Me_{i}=L_{i}\oplus P_{i}$ with $h_{i}(L_{i},L_{i})=0$ and $h_{i}(P_{i},P_{i})=0$ (cf. [8, Section 2.2]). So $M=\bigoplus_{i=1}^{t}L_{i}\oplus P_{i}\cong(\bigoplus_{i=1}^{t}L_{i})\oplus(\bigoplus_{i=1}^{t}P_{i})$. Let $L=\bigoplus_{i=1}^{t}L_{i}$ and $P=\bigoplus_{i=1}^{t}P_{i}$. We check that $h(L,L)=0$; the proof of $h(P,P)=0$ is similar. It suffices to show that $h(\ell_{i},\ell^{\prime}_{j})=0$ for each $\ell_{i}\in L_{i}$ and $\ell^{\prime}_{j}\in L_{j}$. If $i\not=j$ then $h(\ell_{i},\ell^{\prime}_{j})=0$ and if $i=j$ then $h(\ell_{i},\ell^{\prime}_{i})=h_{i}(\ell_{i},\ell^{\prime}_{i})=0$.
Every $\alpha\in\operatorname{Sper}^{\mathrm{max}}R_{i}$ can be seen as an element $\alpha^{\prime}$ in $\operatorname{Sper}^{\mathrm{max}}R$, and a direct verification of the definition of signature shows that $\operatorname{sign}^{\mathscr{M}}_{\alpha^{\prime}}h=\operatorname{sign}^{\mathscr{M}}_{\alpha}h_{i}$. Therefore, (2) gives that $\operatorname{sign}^{\mathscr{M}}_{\alpha}h_{i}=0$ for every $\alpha\in\operatorname{Sper}^{\mathrm{max}}R_{i}$ and every $i=1,\ldots,t$. By Propositions 5.3 and 5.4 we have that for every $i=1,\ldots,t$ there exists $n_{i}\in\mathbb{N}$ such that $2^{n_{i}}\times h_{i}$ is hyperbolic. Thus, letting $n=n_{1}+\cdots+n_{t}$, it follows that $2^{n}\times h_{i}$ is hyperbolic for $i=1,\ldots,t$ and hence that $2^{n}\times h$ is hyperbolic. ∎
Acknowledgements
We thank Igor Klep for many stimulating exchanges of ideas, and are very grateful to the Fondazione Bruno Kessler–Centro Internazionale per la Ricerca Matematica and Augusto Micheletti for hosting the three of us in December 2018 for a Research in Pairs stay during which some of the ideas that eventually led to this paper were discussed.
We are very grateful to the Referees for reading our paper in great detail, for spotting a serious mistake, and for the many pertinent and helpful questions and suggestions, which led to substantial improvements.